### v1

# Conghui He4, Xiaoyang Wang6, Fan Wu1

Abstract
Large language model (LLM) and agent techniques for data analysis (a.k.a LLM/Agent-as-Data-Analyst) have demonstrated substantial impact in both academica and industry. In comparison with traditional rule or small-model based approaches, (agentic) LLMs enable complex data understanding, natural language interfaces, semantic analysis functions, and autonomous pipeline orchestration. The technical evolution further distills five key design goals for intelligent data analysis agents, namely semantic-aware design, modality-hybrid integration, autonomous pipelines, tool-augmented workflows, and support for open-world tasks. From a modality perspective, we review LLM-based techniques for $(i)$ structured data (e.g., table question answering for relational data and NL2GQL for graph data), $(ii)$ semi-structured data (e.g., markup languages understanding and semi-structured table modeling), $(iii)$ unstructured data (e.g., chart understanding, document understanding, programming languages vulnerable detection), and $(iv)$ heterogeneous data (e.g., data retrieval and modality alignment for data lakes). Finally, we outline the remaining challenges and propose several insights and practical directions for advancing LLM/Agent-powered data analysis.
Index Terms:
LLM, Agent, Data Analysis, Structured Data, Semi-Structured Data, Unstructured Data, Heterogeneous Data
Figure 2: Technical Overview of LLM/Agent-as-Data-Analyst.
1 INTRODUCTION
Data analysis, broadly defined as the process of inspecting, transforming, and modeling data to discover useful information and support decision-making, constitutes a cornerstone of modern scientific research and business intelligence hardy2004handbook ; temporal2022 ; LLM4DM ; chat2data2024 . It spans a wide spectrum of data modalities, from structured databases and semi-structured tables to unstructured documents and videos, underpinning critical applications across domains such as finance, healthcare, engineering, and social sciences nooralahzadeh2024explainable ; batko2022use ; cao2022ai .
1.1 Limitations of Conventional Data Analysis
Traditional data analysis pipelines, although effective in extracting information and statistical patterns, often demand substantial domain expertise, manual feature engineering, and the integration of multiple specialized tools xiao2024cellagent . These limitations become more pronounced as data grows in scale, complexity, and heterogeneity liang2024natnl2gql , ultimately constituting the inherent weakness of conventional data analysis.
L1: Manual Development. The most labor-intensive challenge is that traditional data analysis workflows rely on rigid, manually designed pipelines that require domain experts to specify modeling steps in advance zhou2025cracksql ; zhou2025cracksqldemo ; zhang2023reactable . In large-scale enterprise databases, where separate tables for customers, orders, payments, shipments, and products each contain dozens of attributes, answering even simple business queries such as “What is the average delivery time for premium customers in the last quarter?” may necessitate the analysts to choose related tables, find corresponding attributes, process complex multi-table joins, apply domain-specific constraints, and reconcile the timestamp fields. For document query answering, retrieving query-relevant information from long documents containing diverse elements (e.g., text blocks, tables, charts, images) typically requires analysts to examine pages sequentially and summarize the content, which is a highly labor-intensive process. Such handcrafted pipelines are time-consuming to build and difficult to adapt when data characteristics or analysis goals change.
L2: Hard-Coded Tool Dependency.One observation is that existing data analysis systems often come with a suite of tightly coupled built-in tools (e.g., visualization dashboards, statistical packages, or machine learning modules torchptpth ) tailored to specific workflows, which stems from the inherent complexity of the underlying data types. For instance, to extract statistical information from a knowledge graph and subsequently model it with linear regression, a human analyst must first formulate GQL queries and invoke the built-in APIs of graph databases to obtain the data, and then employ machine learning libraries (e.g., scikit-learn, PyTorch) for further analysis. In contrast, analyzing 3D models relies more heavily on domain-specific expertise and specialized systems (e.g., AutoCAD, PyMOL pymol ). Such analyses are often constrained to the functionalities embedded in these systems, requiring analysts, particularly those without coding expertise, to integrate multiple systems in order to complete a single task. The integration of heterogeneous systems not only increases the complexity and overhead of the analysis process, but also hampers extensibility and complicates integration into broader analytical workflows.
L3: Homogeneous Modality Support.Current data analysis systems are generally optimized for specific data types, with architectures and query engines tailored to particular formats. For example, relational databases are well-suited for structured data LLMQO , while systems such as MongoDB are designed for semi-structured formats like JSON mongodb . When analyzing videos with accompanying timestamped descriptions, human analysts must manually align frames with their corresponding documents to enable cross-modal analysis mu2024snagscalableaccuratevideo . This specialization hampers the analysis of heterogeneous data across modalities, thereby increasing analytical complexity, introducing errors, and limiting both the scope and efficiency of data-driven insights.
L4: Basic Format-Based Analysis.While conventional analysis methods efficiently support operations such as filtering, aggregation, they lack the ability to reason over semantic information embedded in data. For example, in table analysis, if a cell contains lengthy textual paragraphs, SQL queries can only retrieve or match keywords without understanding the semantics (e.g., sentiment, intent, causal relationships) zhang2023reactable . Likewise, for unstructured data such as charts or documents, analysts must manually interpret and summarize content to fulfill analytical needs han2023chartllama . The absence of semantic understanding constrains the analytical depth of current systems and limits their ability to address complex, knowledge-intensive tasks that demand reasoning beyond surface-level patterns.
1.2 Opportunities by LLM/Agent-As-Data-Analyst
Recent advances in large language models (LLMs) and LLM-based agents have introduced new opportunities to alleviate these challenges. As shown in Figure 1, by enabling auto-designed analysis pipelines, adaptive tool-assisted workflows, and natural language interaction, LLMs hold the potential to lower technical barriers, enhance interpretability, and accelerate the discovery of actionable insights from diverse forms of data.
O1: Complex Data Understanding. LLMs are capable of processing and reasoning across complex data, including relational data, semi-structured tables, and unstructured texts, owing to their ability to capture latent patterns and contextual dependencies su2024tablegpt2 ; tang2024graphgpt . For example, they can not only comprehend the semantic content associated with nodes and edges but can also reason about the underlying structural properties (e.g., connectivity, community patterns, or hierarchical organizations). Such a deep and holistic understanding empowers analysts to derive richer insights that integrate quantitative measures with qualitative nuances arora2023language .
O2: NL-Based Interface. An LLM-powered analysis agent leverages natural language as the primary interface for interaction, enabling users to articulate analytical requirements without the need for specialized query languages such as SQL or advanced programming skills liang2024natnl2gql ; birdsql . This natural language interface enhances convenience for analysts, accelerates the overall analysis process, and broadens participation in data-driven decision-making across diverse user groups tang2025st ; han2023chartllama .
O3: Semantic Operators. LLMs enable semantic-level operations such as structural interpretation and content summarization, which are particularly beneficial for data rich in structural relationships or textual content wang2024interactive . For example, when querying a medical record database, a conventional query may retrieve only documents containing the term “fever”, whereas an LLM-based operator can also recognize semantically related expressions such as “high temperature” or “febrile condition”. This semantic enrichment enhances the expressive power of queries, enabling analysts to obtain more accurate, context-aware results jiang2025explainable ; wang2024chainoftable .
O4: Autonomous Evolving. LLM agents function not as static tools but as adaptive systems capable of refining their performance through continuous interaction and learning. Such adaptability is essential, as analytical requirements and data environments inevitably evolve over time hong2024datainterpreter . For instance, a financial analysis agent may initially concentrate on trend forecasting but, through iterative feedback, expand its capabilities to detect anomalies in real time or incorporate emerging risk factors. This capacity for autonomous evolution ensures that the system remains advanced, alleviates the need for manual reconfiguration, and delivers increasingly personalized and efficient analytical support atreja2025alasautonomouslearningagent .
1.3 Techniques of LLM-Powered Data Analysis
As shown in Figure 1 and Figure 2, given the diversity of data formats and application scenarios, we establish a taxonomy of LLM-powered data analysis which progresses along two dimensions: (1) the range of data modalities supported, i.e., structured, semi-structured, unstructured, and heterogeneous (the x-axis), and (2) the evolution of interaction paradigms, i.e., code-based, domain-specific language (DSL)-based, and natural language (NL)-based (the y-axis).
LLM/Agent for Structured Data Analysis. Structured data, such as relational databases relationaldata and graphs graphdata , remains central to industrial practice due to its standardized schema and well-defined semantics. Traditional approaches rely on SQL or similar code-based query languages, later extended by DSLs that simplify domain-specific tasks neo4j . With LLMs, users to interact with structured data through NL-interface, further through code generation, DSL mapping, or LLM-based question answering sqlsurvey ; shi2024sqlsurvey , with agents further orchestrating multi-step analytical workflows zhu2024sqlsurvey . The core of LLM driven structured data analysis is to use LLM for pipeline or end-to-end processing.
$\bullet$ Relational Data. Relational data is typically stored and managed through specialized databases. A method for LLM-based relational data analysis is to leverage the LLM as the interface to convert analysis task in natural language into SQL or code. Then manipulate the database through SQL or code to get the analysis results. To better align natural and specialized languages, techniques like schema linking zhang2024finsql , information retrieval DINSQL and task decomposition hong2024datainterpreter are applied to the pipeline; or the model is fine-tuned yin2022code on certain tasks for better end-to-end generation. Another direction is to skip the database manipulation and directly use LLM for semantic analysis. This requires the model to better understand intentions of analysis tasks, and relationships within structured data. Hence RAG patnaik2024cabinet , prompt engineering tatllm and task decomposition zhao2024tapera are used in pipelines for semantic analysis, and MLLM zheng2024multimodal or LLM su2024tablegpt2 are trained or fine-tuned for reasoning based on textual or visual table representations. Time series data, as a specialized form of relational data characterized by temporal dependencies, can also benefit from similar LLM-powered techniques. Recent works alnegheimish2024can ; DBLP:conf/iclr/0005WMCZSCLLPW24 adopt analogous strategies, such as translating natural language into code, retrieving and transforming sequences, or employing LLMs for direct temporal reasoning.
$\bullet$ Graph Data. Graph data represents entities and inter-dependencies to model complex network semantics, presenting challenges due to vast search space and complex path reasoning. There are databases and specified query language for graph data  blazegraph ; graphdb . So LLM can also serve as the interface between natural language and graph query language. Techniques like agent liang2024natnl2gql , fine-tuning and prompt engineering liang2024aligning are introduced for LLMs’ better comprehension of the graph structure and language syntax. And for semantic analysis based on the graph, The mainstream methods are reflect the ideas of RAGzhang2022subgraph , agent luo2025kbqa and fine-tuning ye2024instructglm .
LLM/Agent for Semi-structured Data Analysis. Semi-structured data lies between unstructured text and fully structured relational databases. It typically contains some organizational structure but does not conform to rigid schemas.
$\bullet$ Markup Language. XML, JSON, and HTML are common forms of markup language data. As markup languages consist of both tags and content, and inherently exhibit structural properties, their evolution bears strong similarity to that of semi-structured tables, which has also motivated the development of structure-aware PLMs arora2023language ; wang2022webformer .
$\bullet$ Semi-Structured Tables represent a flexible form of tabular data whose evolution is closely tied to advances in structured table understanding and the semantic capabilities of LLMs. Traditional approaches, often based on PLMs(herzig2020tapas RoBERTa ), struggle to capture the complex layouts, irregular headers, and hierarchical structures due to their limited ability to encode semantics. With the advent of LLMs, new paradigms have emerged for handling such tables, including SemiStru2Stru transformations that convert semi-structured tables into relational-like forms 10.1145/3726302.3730071 ; li2023autotablessynthesizingmultisteptransformations , and DSL-based modeling that leverages domain-specific languages to explicitly encode structure and operations tang2025st .
LLM/Agent for Unstructured Data Analysis. Unstructured data encompass a wide range of modalities such as chart, video, document and 3D models, which lack a fixed schema and therefore pose significant challenges for conventional analysis pipelines.
$\bullet$ Chart. Traditional chart analysis systems typically rely on handcrafted feature extraction, template matching, or rule-based parsing techniques kafle2018dvqaunderstandingdatavisualizations ; mittal-etal-1998-describing ; reiter-2007-architecture , which often struggle with variations in design, layout, and data representation. With the advent of LLMs, we can leverage multimodal understanding capabilities to interpret visual and structural elements of chartsmasry2023unichart , generate contextual descriptions, perform semantic data extraction, and support natural language-based reasoning masry2023unichart , captioning liu2024chartthinker ; obeid-hoque-2020-chart ; singh2025figcapshffiguretocaptiongenerativeframework and QA wu2024chartinsights ; das2025chartsofthoughtenhancingllmvisualization ; xu2025chartmoemixturediverselyaligned tasks over different graphical data.
$\bullet$ Video. Video encodes spatial content that evolves over time, requiring joint modeling of semantics and dynamics. Conventional methods rely on vision backbones with temporal pooling or attention, which struggle with annotation costs and long-sequence efficiency chen2024timemarker wang2024grounded . Recent advances reformulate videos as structured token sequences for LLM reasoning. Agents further decompose queries into temporal grounding, multimodal fusion, and summarization, enabling richer temporal understanding and efficient computation chen2024timemarker deng2025seq2time . Building upon this, LLMs have also been extended to Video Emotional Analysis , where multimodal fusion of visual, acoustic, and textual cues enables affective state inference.In addition, posture-based emotion analysis and 3D mesh reconstruction support nuanced social interaction modeling and relational emotional predictions.jadhav2023ai ; muller2024predicting . For Object Detection, integrating high-precision detectors with multimodal reasoning allows object-centric summarization and reference grounding in videos to make correct object detections in videos de2024video ; yuan2025videorefer . Furthermore, Gesture and Behavior Detection leverages LLM-driven pipelines extract fine-grained motion features and gestures, supporting interaction analysis and embodied behavior reasoning, often via 3D reconstructions for higher fidelity whitehead2025utilizing ; weng2025artificial .
$\bullet$ Document. Documents, such as PDFs, web pages, and scanned reports, serve as primary carriers of information in business and academia. Traditional document analysis relies on Optical Character Recognition (OCR) and rule-based template matching dococr ; ha2022information , which are often ineffective for complex or varied layouts and fail to grasp the deep semantics of the content. LLMs, particularly Multimodal LLMs, have revolutionized document understanding by unifying the processing of text, layout, and visual information in three aspects: (1) Architectural innovations from the LayoutLM series to DocLLM, enable a synergistic understanding of document structure and content xu2020layoutlm ; xu2021layoutlmv2 ; huang2022layoutlmv3 ; wang2024docllm . Concurrently, (2) Retrieval-Augmented Generation (RAG) is employed for question answering and summarization across long or multiple documents lewis2020retrieval ; lei2024visdommultimodal , while (3) synthetic data generation effectively alleviates the scarcity of labeled data seol2024posterllama ; shi2025layoutcot . These techniques collectively drive the shift from simple information extraction to deep document reasoning and synthesis.
$\bullet$ Program. Program analysis, which aims to understand, verify, and optimize source code, is a cornerstone of software engineering. Traditional methods depend on static analysis (e.g., compiler theory) and dynamic analysis (e.g., unit testing), which, while rigorous, require extensive expertise and are difficult to scale to large, semantically complex codebases. By learning from vast amounts of code, LLMs have acquired powerful capabilities for code generation and comprehension, giving rise to new analytical paradigms. The core of this advancement lies in the construction and utilization of code-task pairs—paired data of code snippets with their specific attributes (e.g., vulnerability labels, functional descriptions). Through advanced data synthesis techniques like iterative refinement and self-correction luo2023wizardcoder ; chen2023teaching , models can generate high-quality training data, thereby empowering frontier applications such as Automated Theorem Proving (ATP), vulnerability detection, and repository-level code completion xin2024deepseek-prover ; liu2023gptvulnerability ; zhang2023repocoder . Notably, the deep evolution of Retrieval-Augmented Generation (RAG) in the code domain has significantly enhanced the model’s ability to leverage context from entire repositories liang2024repofuse ; wu2024repoformer ; guo2024ft2ra .
$\bullet$ 3D Model. 3D models represent objects or scenes in Euclidean space, expressed as point clouds, meshes, or voxels, and are widely applied in scene understanding and scientific analysis. Conventional pipelines rely on geometric processing (e.g., mesh simplification, point cloud registration) hoppe1997view ; garland1997surface ; chen1992object and modeling software such as Blender or Maya blain2019complete ; tickoo2018autodesk , which require manual annotation and lack semantic understanding. Recent LLM-based approaches enable 3D–language alignment hong20233d , where geometry is transformed into structured embeddings or textual descriptions for reasoning. Agents orchestrate specialized 3D encoders and toolchains to support downstream tasks such as captioning, navigation, and scientific QA hong20233d ; xiong20253ur . Building on this foundation, 3D-Language Fusion provides unified frameworks that map point clouds and meshes into embeddings aligned with natural language, enabling tasks such as captioning and QA 3D-LLM hong20233d , 3UR-LLM xiong20253ur Domain-specific extensions integrate molecular and protein structures into multimodal reasoning, as in 3D-MoLM li2024towards , ProteinChat guo2023proteinchat , and ProtChatGPT wang2024protchatgpt . Moreover, 3D-Derived Task Enhancement leverages multi-agent systems leverage textual mediation of geometry for scene description, navigation, and retrieval, improving interpretability and efficiency hong20233d ; xiong20253ur ; li2024towards . Finally,Cross-Modal Refinement employs feature enhancement and domain adaptation techniques bridge 2D and 3D modalities, strengthening generalization in multimodal LLMs (e.g., visual grounding, 2D–3D alignment, or joint pretraining) hong20233d ; li2024towards ; li2024towards ; xiong20253ur .
LLM/Agent for Heterogeneous Data Analysis. Heterogeneous data refers to the integration of diverse data types (e.g., relational data, semi-structured tables, document images) wang2017heterogeneous . Early research focused on heterogeneous data management iceberg , which supports data retrieval through SQL-like languages. More recent advances with LLMs address three main directions: (i) modality alignment across data types Unicorn ; chen2023symphony (e.g., leveraging natural language descriptions to compute cross-modal similarity), (ii) natural language interfaces for heterogeneous data retrieval patel2024lotus ; wang2024must (e.g., translating user queries into sequences of predefined APIs), and (iii) heterogeneous data analysis agents nooralahzadeh2024explainable ; wang2024interactive (e.g., equipping LLMs with semantic operation tools tailored to different modalities).
LLM/Agent for Data Analysis Evolution. As shown in Figure 1, the technological evolution of LLM-powered Data Analysis Agents can be summarized along five key dimensions, each corresponding to the design goals of a unified data analysis agent.
$\bullet$ From Literal to Semantic. Early LLM-powered data analysis agents primarily assist data analysis by generating executable code or domain-specific languages (DSLs). Their evolution has shifted toward leveraging semantic understanding to derive analytical results directly, reducing the need for intermediary procedural steps.
$\bullet$ From Modality-Specific to Modality-Hybrid. Initial data analysis agents are limited to analyzing a single type of data. The current trend emphasizes the integration of multiple data modalities, enabling agents to perform coordinated analyses across heterogeneous data.
$\bullet$ From Manual Development to Autonomous Design. Early agents require manual decomposition of analytical workflows (e.g., task decomposition, code generation, operation execution). Modern agents increasingly support autonomous workflow design, with broader operational capabilities and reduced human intervention.
$\bullet$ From Tool-Coupled to Tool-Assisted. Traditional agents rely on tightly integrated, framework-specific tools. The evolution favors decoupled architectures, where LLMs can leverage arbitrary toolsets, enhancing flexibility and adaptability in diverse analytical contexts.
$\bullet$ From Closed World to Open World. Initially, agents are predominantly tailored to domain-specific tasks (e.g., financial or industry analysis). The trend is toward general-purpose agents capable of analyzing diverse and real-world data (e.g., documents, videos), lowering user entry barriers.
1.4 Comparison & Contributions
Compared with existing surveys on LLMs for data analysis sqlsurvey ; shi2024sqlsurvey ; zhu2024sqlsurvey ; shang2024graphsurvey ; ren2024graphsurvey ; jin2022tablesurvey ; wu2025tablesurvey ; ding2024docsurvey ; barboule2025docsurvey ; farahani2023chartsurvey ; al2024chartsurvey ; tang2025videosurvey , our work provides a more comprehensive and detailed overview of the key techniques applied across different data types, while also emphasizing the interrelationships among these data types. We uniquely examine the development trends from the perspectives of data modalities and interface languages, and further outline key dimensions for the design of a general LLM-based data analysis agent.
$\bullet$ Full-Vision Introduction to Typical Data Analysis Tasks. Different from existing surveys that typically focus on a single modality or task (e.g., NL2SQL sqlsurvey ; shi2024sqlsurvey ; zhu2024sqlsurvey , graph understanding shang2024graphsurvey ; ren2024graphsurvey , table question answering jin2022tablesurvey ; wu2025tablesurvey , document understanding ding2024docsurvey ; barboule2025docsurvey , chart understanding farahani2023chartsurvey ; al2024chartsurvey , video understanding tang2025videosurvey ), we systematically organize the landscape of data analysis by categorizing techniques across structured, semi-structured, unstructured, and heterogeneous data. This roadmap also traces the technical evolution of LLM-powered data analysis and identifies five key dimensions that serve as design goals for a general-purpose data analysis agent.
$\bullet$ Detailed Review of Data Analysis Techniques. Beyond high-level summaries, we provide an in-depth examination of representative methods, discussing their underlying principles, technical designs, and application scenarios. In contrast to existing surveys, we further emphasize the critical role of data curation methods tailored to specific downstream tasks and offer corresponding data analysis insights.
$\bullet$ Recent Advances in LLM/Agent-as-Data-Analyst. In addition to established techniques, we emphasize the latest developments that leverage LLM for data analysis (e.g., agentic design, multimodal alignment, interaction techniques). By incorporating these cutting-edge advances, our survey provides an up-to-date reference for researchers and practitioners seeking to understand the state-of-the-art.
$\bullet$ Open Challenges and Future Directions. We identify the key technical and practical challenges that remain unresolved, such as scalability, evaluation, and integration into real-world systems. Building on these insights, we also outline promising future research directions to guide the development of general-purpose LLM-based data analysis agents.
1.5 Organization of Our Survey
Section 2 discusses LLMs for structured data analysis, covering relational data (Section 2.1) and graph data (Section 2.2). Section 3 reviews LLMs for semi-structured data analysis, including markup languages (Section 3.1) and semi-structured tables (Section 3.2). Section 4 examines LLMs for unstructured data analysis, encompassing charts (Section 4.1), videos (Section 4.2), documents (Section 4.3), programming languages (Section 4.4), and 3D models (Section 4.5). Section 5 addresses LLMs for heterogeneous data analysis. For each data type, we first present the data analysis techniques, followed by a subsection on data curation. Finally, we discuss the challenges and future directions associated with each data type in Section 6, and conclude the survey in Section 7.
2 LLM for Structured Data Analysis
Takeaway: For relational data, LLM takes part in natural language interfaces, semantic analysis and time series analysis. For graph data, LLM can perform graph analysis query generation and semantic analysis.
Structured data refers to data with well-defined schemas like relational data relationaldata and graph data graphdata . The commonality of them is pre-defined patterns that ensure clear organization and efficient querying. Relational data is characterized by row and column tables, or associated keys. Graph data models entities and their relationships through nodes and edges.
2.1 Relational Data Analysis
LLM for Natural Language Interfaces. Basic analysis jobs for relational data are typically characterized by well-defined operations. These operations include basic calculations (e.g., summation, averaging, counting, ranking) relationaldata , statistical analysis (e.g., regression and clustering) hoff2015multilinear ; taskar2001probabilistic , and data quality assurance processes (e.g., constraint validation, outlier detection) maervoet2012outlier . Such tasks can generally be supported by tools like SQL or Python libraries (e.g., Pandas). Traditional methods typically rely on strictly defined SQL language and pre-defined analysis rules or algorithms. In contrast, LLMs can not only perform conventional analysis tasks, but also enable natural language–to–SQL/code conversion and advanced semantic analysis, leveraging strong language understanding and reasoning capabilities.
Figure 3: LLM for Structured Data Workflows - (a) Pipeline Method. (b) End-to-End Method.
$\bullet$ NL2SQL. With the help of LLM, users can directly perform operations using natural language. NL2SQL focuses on translating natural language queries into SQL commands by leveraging techniques such as $(i)$ schema linking, which aligns user intents with database schema to resolve ambiguities and filters out irrelevant schema to enhance the conciseness of input zhang2024finsql ; petsql , $(ii)$ content retrieval, which dynamically extracts relevant information from the database to refine query generation talaei2024chess ; li2024codes , and $(iii)$ SQL generation strategies such as multi-step generation that concentrates on different SQL parts at each step fan2024combining ; zhou2025cracksql ; zhou2025cracksqldemo , intermediate SQL representation or hybrid table representation for better LLM comprehension DINSQL ; DBLP:journals/pacmmod/XieXZG25 ; long2025bridging , and different decoding strategies (e.g., beam search, greedy search, or PICARD SQL360 ). Other surveys provide more comprehensive overviews for NL2SQL techniques Liu2024ASO ; luo2025natural ; shi2024survey .
$\bullet$ NL2Code. Different from NL2SQL, NL2Code approaches emphasize enhancing relational data analysis through generating Python code (e.g., Pandas, NumPy), which includes a vast number of library APIs characterized by high variability and complexity, and often requiring the handling of complex chain operations. Recent studies have made progress in alleviating these challenges, though many remain open.
(1) Model Fine-Tuning: PACHINCO yin2022code fine-tunes a 62B parameter PALM chowdhery2022palm model in two stages (i.e., separately using a Python source code corpus with 64B tokens and a Jupyter notebook corpus with 9.6B tokens) so as to improve model performance on analysis-related tasks (e.g., calculate the amount of games added in each year for each month). DataCoder huang2024datacoder utilizes different types of contexts (e.g., code, text, and data) by employing dual encoders (e.g., data encoder and code + text encoder) and one general decoder to generate code in notebooks.
(2) LLM-Based Analysis Agent: Data Interpreter hong2024datainterpreter leverages LLMs through APIs to generate task and action graphs. Specifically, they utilize LLM’s semantic reasoning ability to accurately decompose complex user queries into subproblems (e.g., correlation analysis, data exploration, and anomaly detection), and further refine and verify each subproblem to improve code generation results for data science tasks. Another approach bai2025collaboration fine-tunes a BART model DBLP:conf/acl/LewisLGGMLSZ20 as an intelligent agent with datasets constructed from task-instruction and role-definition prompts. The agent further generates more detailed prompts to guide LLMs to complete the NL2Code task step-by-step (e.g., role definition, demand optimization, code writing, and code review).
LLM for Semantic Analysis. Moreover, some jobs require LLM-based analysis, such as those that involve semantic understanding or demand outputs in natural language format (e.g., table summarization). These challenges call for methodologies like (1) multi-step question answering (QA) with diverse decomposition strategies and (2) end-to-end QA leveraging specifically optimized LLMs.
$\bullet$ Multi-Step QA. Multi-step question answering (QA) refers to decomposing complex queries into a sequence of sub-questions to facilitate step-by-step reasoning. According to the question decomposition mechanisms, existing methods can be categorized into two types: (1) static decomposition, which follows pre-defined and fixed processing steps (e.g., retrieve-select-reason), and (2) LLM-driven iterative decomposition, in which the LLM dynamically determines the next operation based on the contextual history of the reasoning process.
(1) Static Decomposition: The static decomposition includes Retriever-Selector-Reasoner frameworks and the variants, which partition tasks into modular components for better multi-step inference and enhanced interpretability. The Extractor-Reasoner-Executor paradigm tatllm extracts the relevant segments from the context, generates the logic rules or equations, and performs the rules or executes the equations to get the final answer through LLM prompting. S3HQA lei2023s3hqa employs a retriever for initial filtering of heterogeneous resources, a selector for identifying the most relevant factual knowledge, and a generation-based reasoner for producing the final answer. For multi-table scenarios, SGAM wang2025plugging encodes schema links and join paths as a graph, transforms the query into interpretable reasoning chains on the graph, and executes along the chains to obtain the final answer.
(2) Iterative Decomposition: Static decomposition paradigm often struggles with multi-hop queries, while LLM-driven iterative decomposition overcomes these limitations by dynamically refining sub-tasks through recursive reasoning.
TAPERA zhao2024tapera introduces the query decomposition step into the question answering process by adopting the LLM-driven approach. A Planner first decomposes the query into sub-queries, forming an initial plan. The Reasoner then generates executable programs for each sub-query, and the Answer Generator derives answers based on the program outputs to fulfill the plan. Finally, the Planner updates or finalizes the plan as needed.
Similarly, ReAcTable zhang2023reactable and CHAIN-OF-TABLE wang2024chainoftable iteratively generate operations and update the table to present a reasoning chain as a proxy for intermediate thoughts through prompting LLMs and in-context learning.
$\bullet$ End-to-End QA. End-to-End Question Answering (QA) refers to approaches in which the answer-generating LLM directly produces the final response without intermediate steps or iterative refinement. Based on the data representation and processing mechanisms, the relevant methods can be classified into table-specific LLM fine-tuning, table content retrieval, and table-as-image analysis.
(1) Table-Specific LLM Fine-Tuning: Fine-tuning LLMs on task-specific table datasets enables them to internalize analytical knowledge directly within their parameters. TableGPT li2023tablegpt fine-tunes LLMs like GPT-3.5 using a diverse set of table tasks synthesized from real-world tables. Building on Qwen2.5 qwen2025qwen25 , TableGPT2 su2024tablegpt2 introduces a table encoder to generate a hybrid table representation, an adapter to generate query representations, and a LLM decoder to generate an agent workflow (i.e., the tool execution pipeline) for deriving the final answer. The TableGPT2 model is pre-trained on 593.8K tables and fine-tuned 2.36M question-answer pairs.
(2) Table Content Retrieval: Instead of embedding the whole table, table content retrieval enhances model performance by eliminating noisy parts of the table while retaining information relevant to question answering. CABINET patnaik2024cabinet employs a weakly supervised component to produce a parsing statement that defines the criteria for selecting relevant rows and columns, emphasizing the corresponding table cell content. TableMaster cao2025tablemaster constructs a refined subtable through row and column lookup. By leveraging carefully designed LLM prompts (e.g., provide objective, table definition, table information, question, instructions, and response format), it ranks all candidate columns, selects a relevant subset based on the query, and then instructs the LLM to generate an SQL query for extracting the most relevant rows. In the test of Multi-table and Multi-hop Question Answering dataset (MMQA) wu2025mmqa , LLMs are required to retrieve tables from a given corpus and then perform multiple inferences. A multi-table retrieval method (MTR) that iteratively retrieves question-related and table-related tables achieves the best performance in the test.
(3) Table-As-Image Analysis: Due to the limitations of (text-only) LLMs in understanding table structures, the table-as-image approach has been proposed, converting tables into images for analysis using multimodal LLMs. Table-LLaVA zheng2024multimodal applies incremental pretraining to LLaVA-7B liu2024llava on 150K table recognition samples (e.g., input a table image and output table representations in HTML, Markdown, or LaTeX), enabling the model to align table structures and elements with textual modality. It is further fine-tuned on 232K samples on question answering, text generation, fact verification, and structure understanding tasks to enhance its instruction-following ability. To enable a single model to perform various analytical tasks, TabPedia zhao2024tabpedia introduces the concept synergy mechanism, abstracting all table analysis tasks into concepts. Built on Vicuna-7B zheng2023judging , it appends meditative tokens to the input of the LLM decoder, which adaptively activates different regions of visual tokens and helps the model interpret the intent behind specific task questions. However, such methods face limitations when processing twisted or distorted tables, and their performance degrades significantly when directly handling document images.
LLM for Time Series Analysis. Time series data is a special type of relational data that can be stored in relational databases or in specialized time series databases like influxdb naqvi2017time . It represents the sequential data obtained by observing real or simulated scenes in chronological order, with each time step corresponding to one or more variables. Time series analysis typically aims to learn the trend or periodic pattern of data changes over time (e.g., sensor data, price fluctuation, traffic flow, etc.), and to model interdependence among variables (e.g., causality, correlation, coupling), enabling tasks such as prediction, classification, and detection liu2025towards . If viewed as a wide table indexed by time, time series analysis emphasizes the dynamic evolution of parameters along the timeline, in contrast to relational data analysis, which focuses on static relationships.
Traditional time series analysis includes statistical-based and machine-learning-based methods, both rely on certain models such as ARIMA and LSTM 8614252 . However, statistical methods have shortcomings such as assuming linear relationships, tedious manual feature engineering, and limited ability for complex nonlinear patterns add rahmani2022association . The development of machine learning and deep learning provides a new solution for time series analysis 9005997 . At the same time, people are trying to improve the interpretability of analysis result 9713986 . Recently, LLMs for time series analysis have become a new focus due to their sequence modeling capabilities, less data dependencies, and the potential to capture long-term dependencies.
$\bullet$ TS2NL. Due to the shared sequential features of time series data and natural language, the TS2NL method directly converts time series data into natural language and provides it to LLMs. The format of natural language brings better intuitiveness and interpretability.
(1) End-to-End Reasoning: End-to-end reasoning requires the answer-generating LLM to directly generate the result with the provided prompt, which contains the time series data, instruction, and other context (e.g., news, events and logs). For example, SIGLLM alnegheimish2024can studies two paradigms for time series anomaly detection with LLMs. It mainly studies two paradigms: directly requesting LLMs for detection, or requesting them for prediction and then comparing the predicted results with actual data. TimeRAG yang2025timerag utilizes similar historical time series data to help the large model understand the current scene, achieving domain adaptation without a lot of training. It constructs a knowledge base with historical time series data, and retrieves reference sequence from it to strengthen the prediction. The reference sequences are selected by the distance measured by dynamic time warping (DTW) senin2008dynamic . Some work also studies the zero-shot or few-shot forecasting capability of LLMs in specific fields such as finance yu2023temporal and climate DBLP:journals/corr/abs-2411-13724 .
(2) Agent-Based Methods: Agent-based methods utilize LLM agents to collaborate to complete analysis. Different agents often complete their own tasks under respective prompts. TimeCAP lee2025timecap is a framework with two agents and a multi-modal encoder. The summary agent generates a contextual summary of time series data, and the prediction agent uses the summary for prediction. The encoder generates predictions using both the data and summary. It also samples relevant text to reinforce the prediction agent. The final prediction is derived from the linear combination of the encoder prediction and the agent prediction. TimeXL jiang2025explainable proposes a more complex framework consisting of a multi-modal prototype-based encoder and three GPT-4o agents. The prototypes demonstrate the typical text patterns and temporal patterns. The encoder generates preliminary predictions and explanations based on time series and textual inputs. Three agents are respectively responsible for generating the prediction according to text input and explanations, comparing predicted and actual series to generate the feedback, improving text quality and triggering encoder retraining (e.g., updating the text input of the encoder based on feedback). The final prediction is derived from the weighted average of the encoder prediction and the agent prediction. The final prediction is also derived from the linear combination of the encoder prediction and the agent prediction. Considering the augmentation of prediction by relevant context, wang2024news selects relevant news to enhance forecasting by agents. A reasoning agent is instructed to select news as context of the prediction prompt, and a evaluation agent critiques the selected news and prediction result against the ground truth to update the logic of reasoning agent.
$\bullet$ Time Series Alignment. Because of the modality differences between time series data and natural language, some methods consider aligning the modalities to better utilize LLMs to analyze time series data. These methods solves the challenge that LLMs may overlook temporal patterns, enabling them to better understand the time series format and analyze the dependencies. According to the mainly optimized architecture of models, they can be classified into designed encoder/decoder and fine-tuned LLM body. Their techniques are compared in the table I.
(1) Designed Encoder/Decoder: By designing and training special encoder or decoder modules, time series data is adapted to the input or output formats of LLMs with the backbone of LLMs unchanged. TIME-LLM DBLP:conf/iclr/0005WMCZSCLLPW24 proposes an effective embedding method to integrate time series data and textual context. It trains a patch reprogram module to encode time series data with text prototypes obtained by linearly probing the pre-trained word embeddings, and uses a frozen LLM embedder of Llama-7B to encode prompts as additional prefix. After the frozen LLM generates output embeddings, a trainable projection module is used to get the final forecasts.
To bridge the gap between structure, numerical features and language semantics, SEED li2025seed develops an encoding method with a token-aware structural encoder, a patch projection and alignment module and a semantic reprogramming module. It finally concatenates the tokens of multivariate time series data and task prompt as the input of LLM. Differently, TimeCMA liu2025timecma doesn’t put all the embeddings into LLM for the shortcoming of LLM-based solutions in learning disentangled embeddings. Instead, it encodes time series data into two branches, time series branch and LLM-empowered branch, combining them for multivariate time series forecasting via cross-modality alignment.
(2) Fine-Tuned LLM Body: Based on designed encoder or decoder modules, some methods fine-tune LLM body on time-series tasks to adapt to the modalities of time series data, which further improves the LLMs’ ability to understand and analyze complex time series. CALF liu2025calf focuses on three aspects: embedding tokens, hidden features and prediction results. Accordingly, it applies the cross-modal match module that derives aligned text tokens queried by time tokens, the feature regularization loss between the textual and temporal hidden feature, and the output consistency loss between textual and temporal prediction. They reduce the distribution discrepancy in the corresponding aspect during the training process. S2IP-LLM pan2024sipllm , on the other hand, decomposes time series by seasonal-trend to explicitly encode each component. Tokens of components and relevant prompts are concatenated as input, while the component predictions are combined as the overall prediction. LLM4TS chang2025llm4ts proposes a two-stage fine-tuning strategy: The time series alignment stage aligns LLMs with time series data, while the forecasting fine-tuning stage fine-tunes LLMs on downstream forecasting tasks. In addition, a two-layer aggregation method integrates data of different time scales by pooling the sum of their embeddings, enhancing LLMs ’ ability to understand temporal information with specific time scale. Apart from prediction tasks, LLMFew chen2025large studies LLMs for multivariate time series classification. It introduces a patch-wise temporal convolution encoder (PTCEnc) with stacked causal convolution blocks for the alignment of time series representations and LLM’s input. And LoRA is used to fine-tune LLMs to improve their feature representation learning.
TABLE I: Comparison of Time-Series Analysis Methods.
Method Encoder Backbone Decoder
TIME-LLM Concatenation Linear Projection
SEED Concatenation Linear Projection
TimeCMA Dual-Modality Multivariate Transformer
CALF Dual-Modality LoRA Dual-Modality
S2IP-LLM Decomposition PEFT Combination
LLM4TS Multi-Scale Two-Stage PEFT RevIN
LLMFew PTCEnc LoRA Classification Head
2.2 Graph Data Analysis
Different from relational data, graph data captures entities (vertices) and their interdependencies (edges) to explicitly model complex network semantics (e.g., social networks and knowledge graphs) beyond rigid tabular schemas. This structure is expressive, however, poses unique challenges, notably the vast search space and complex path reasoning involved in multi-hop queries defgraphdata . Compared with relational data analysis, graph data analysis involves more complex jobs like summarization based on the multi-hop relations across the graph vertices and reasoning over text-attributed graphs whose nodes and edges are associated with text liang2024natnl2gql ; zhou2024r3 . Graph data can not only be stored in relational databases, but also be stored and queried in knowledge graphs and accessed through SPARQL in RDF databases (e.g., Blazegraph blazegraph and GraphDB graphdb ) or Cypher in Neo4j neo4j .
Traditional graph analysis (e.g., statistical methods, graph neural network (GNN) based methods) encompasses a spectrum of tasks, including node classification (e.g., categorizing academic papers into research domains), graph classification (e.g., predicting node properties over molecular graphs), link prediction (i.e., inferring latent relationships between graph nodes), community detection (i.e., identifying densely connected subgraphs), anomaly detection (i.e., identifying deviations from expected patterns), graph clustering, and etc. However, these methods have their own limitations. Statistics-based methods fail to handle complex semantic information (e.g., query can be extremely complex and requires human expertise), while graph neural networks (GNNs) exhibit limited generalization capabilities, necessitating task-specific retraining on different tasks.
In contrast, the advent of LLMs offers transformative potential by leveraging their advanced reasoning capacities and cross-domain generalization abilities, which can (1) simplify the query writing costs (e.g., NL interfaces) and (2) achieve semantic-aware analysis unsupported in traditional ones.
Natural Language to Graph Analysis Query. Different from NL2SQL, the syntax of graph query language generation is more complex (i.e., MATCH, LOOKUP, GET and other operations unique to graph data manipulation) and there exist two operation objects (i.e., vertex and edge) zhou2024r3 . By integrating natural language interfaces with graph data, LLMs facilitate flexible and efficient query generation without the need for specialized model architectures.
To enhance LLMs’ comprehension of the complex syntax of Graph Query Language (GQL), $R^{3}$-NL2GQL zhou2024r3 proposes a hybrid approach leveraging relatively small LLM (e.g., LLaMA3-7B) as a selector and GQL rewriter, while employing a larger LLM (e.g., GPT-4) as a reasoner. The selector identifies the necessary CRUD functions, clauses, and schema, while the rewriter refines the query by aligning it with the relevant graph data retrieved by minimum edit distance and semantic similarity calculation. The LLM then synthesizes the aligned question, selected operations, and schema to generate the final GQL query.
To address the limitations of LLMs in planning and collaborating with other LLMs, NAT-NL2GQL liang2024natnl2gql introduces a three-agent framework. The Preprocessor agent constructs context information, including query rewriting, path linking, and the extraction of query-relevant schemas. The Generator agent, an LLM fine-tuned with NL-GQL data, generates GQL statements based on the rewritten queries and extracted schemas. The Refiner agent iteratively enhances the GQL or contextual information by leveraging error feedback from GQL execution results.
To align LLMs with graph databases in specific fields such as finance and medicine,  liang2024aligning proposes a pipleline combining the fine-tuning and the prompt method. NL-GQL pairs are generated by LLMs for fine-tuning, and the consistency is ensured by the background graph databases and two mutual verification self-instruct methods. In addition, due to the importance of relevant schema in the queries generating of LLMs, it will be extracted and integrated into prompts.
Note that, within the context of AI for Science (AI4Science), the integration of LLMs with graph data analysis has also shown significant potential and wide-ranging applications (e.g., treat polymers as graphs and predict their properties li2024graph ; pei2024bio ), which is not the primary focus of this survey.
LLM-Based Semantic Analysis. Furthermore, certain jobs necessitate semantic-aware analysis, such as summarizing textual paragraphs embedded within graph nodes. Based on the adopted LLM strategies, we classify the relevant methods into retrieval-then-reasoning methods, execution-then-reasoning methods, graph-task-based fine-tuning methods, and agent-based methods.
$\bullet$ Retrieval-Then-Reasoning. Retrieval-then-reasoning first extracts a question-specific subgraph from the graph to identify the most relevant entities and then generates answers using LLMs. To address the challenge of a vast search space, zhang2022subgraph introduces a two-stage approach. First, a trainable and decoupled subgraph retriever selects a relevant subgraph based on the query. Then, reasoning is performed over the retrieved subgraph to derive the final answer. UniKGQA jiang2023unikgqa integrates retrieval and reasoning within a unified model architecture. It consists of a semantic matching module and an information propagation module where the former employs a pre-trained RoBERTa RoBERTa to align questions with graph relations and the latter propagates matching signals along directed edges. Regarding the real-world textual graphs QA, G-Retriever he2024g retrieves relevant nodes and edges based on text encoding, and then uses PCST optimization algorithm to construct subgraphs. The subgraph structure is encoded through Graph Attention Network (GAT) while its content is extracted in text format. These representations are then fed into LLMs to generate the final answer.
$\bullet$ Execution-Then-Reasoning. Execution-then-reasoning refers to the process of parsing natural language queries into executable logical forms (e.g., SPARQL) that align with the graph data, followed by reasoning based on the output of the executed program. Interactive-KBQA xiong2024interactivekbqa introduces an interactive LLM QA framework with a unified SPARQL-based toolset (e.g., entity search, graph pattern search, SPARQL execution, etc.) designed to address complex queries. MCTS-KBQA xiong2025mcts introduces the MCTS method into KBQA, which iteratively executes selection, expansion, evaluation, backpropagation and termination in the graph. A step-wise reward mechanism was designed, using LLMs to evaluate and score each step. FlexKBQA li2024flexkbqa addresses the challenge of lacking high-quality annotated data in real-world scenarios. By prompting LLMs as program translators, it samples program-answer pairs from the knowledge base and generates corresponding natural language questions. The synthetic question-program-answer dataset is used to train lightweight models through execution-guided self-training, which are subsequently employed to annotate real user queries. This approach addresses the distribution shifts between synthetic and actual data, leading to significant improvements in few-shot learning scenarios.
$\bullet$ Graph-Task-Based Fine-Tuning Methods. InstructGLM ye2024instructglm enables generative graph learning by fine-tuning an LLM and leveraging natural language descriptions of graph structures (e.g., offer the first node and the 1-/2-/3-hop neighbors’ information). InstructGraph wang2024instructgraph introduces a stricter code-like graph representation format which constructs entities and triples in the form of list, whose backbone LLM (LLaMA2-7B) is fine-tuned on a graph-centric corpus comprising 1.6 million instances. To mitigate the issue of hallucination, it incorporates Direct Preference Optimization (DPO) algorithm rafailov2024dpo for preference alignment. GraphGPT tang2024graphgpt enhances model performance in zero-shot scenarios by incorporating a structural information encoding module based on Graph-SAGE 10.5555/3294771.3294869 and GCN kipf2017gcn . It fine-tunes the projector bridging the graph encoder and the LLM decoder to align the language capabilities of the foundation LLM (Vicuna-7B) with the graph learning tasks. GLaM dernbach2024glam fine-tunes LLMs to integrate domain-specific knowledge graphs directly into them, which enhances their reasoning capacity. It iteratively partitions and encodes the neighborhood subgraph around each node in a knowledge graph to obtain the context and QA data for fine-tuning.
$\bullet$ Agent-Based Methods. Agent-based methods involve leveraging LLM-based agents with predefined tools (e.g., human-written interfaces or graph processing library APIs) that iteratively interact with the graph data to retrieve, refine, and operate information. StructGPT jiang2023structgpt introduces an iterative reading-then-reasoning framework, leveraging specialized interfaces to operate on graph data. It repeatedly applies an invoke-linearize-generate procedure to derive query results. KBQA-o1 luo2025kbqa utilizes a ReAct agent and MCTS to explore the knowledge base. During the exploration process, the agent gradually generates the logical form of knowledge base operations. The MCTS with policy and reward models helps balance the exploration’s performance and search space. Another approach is to generate an entire reasoning path based on the query and refine it only when necessary. Readi cheng2024necessary initially constructs a reasoning path and instantiates it on the graph. When execution errors occur, it collects error messages and invokes an LLM to revise the path. The final answer is inferred from the instantiated graphs.
2.3 Structured Data for LLM
Datasets for structured data, consisting of structured inputs, task definitions, and labels, play a crucial role in enhancing LLMs for structured data analysis. On one hand, training or fine-tuning with the training set is essential to adapt LLMs to domain-specific knowledge (e.g., finance, medical, climate) and structured data patterns (e.g., relational, time series, and graph). On the other hand, the test sets are widely used as benchmarks to evaluate the performance of LLMs.
The traditional method of dataset construction is usually the combination of real data collection and manual annotation. For example, the widely used table QA dataset WikiTQ pasupat2015compositional uses tables from Wikipedia and question-label pairs annotated by humans. Similarly, the NL2SQL dataset Spider yu2018spider uses databases collected from existing resources and question-SQL pairs annotated by college students. The correctness, diversity and readability of annotation results are ensured through multiple rounds of manual inspection. In addition, the correctness of SQL is further verified by script.
However, for some fields, the data may be scattered in small batches, or kept confidential for business or personal privacy needs, making manual collection difficult. To make matters worse, some manual annotation work not only requires a lot of manpower and time costs, but also requires participants to have a high level of professional knowledge. In this situation, semi-automatic or fully automated dataset generation is increasingly needed.
Relational Data Generation. REaLTabFormer solatorio2023realtabformer can generate relational datasets with the transformer. It first generates the parent table with a GPT-2 model and then generates the relational dataset conditioned on that with a Seq2Seq model. The GPT-2 and Seq2Seq model are trained on their corresponding task respectively. Another method hudovernik2024relational utilizes the structural representation capability of GNNs and the generative capability of diffusion models. It represents a relational database by a heterogeneous graph with the method in  xu2022synthetic , where a parent-child table pair connected by a foreign key is modeled as an attributed bipartite graph. Based on the GNN embeddings of the graph, the diffusion-based TabSyn  zhang2023mixed is used for conditional tabular data generation.
Regarding data augmentation, SyntaxSQLNet DBLP:conf/emnlp/YuYYZWLR18 proposes a method to expand complex SQL datasets. It extracts some universal templates of question-SQL pairs from Spider, and then fill the templates with the schema and values from the target database. Furthermore, CodeS li2024codes achieves domain adaptation with low annotation costs by Bi-directional augmentation. The Question-to-SQL augmentation gives LLMs a few annotated question-SQL pairs, and LLMs can generate new pairs by simulating them. The SQL-to-question augmentation fills in question-SQL templates with the target domain database, and the LLMs then rewrite the question to more natural language.
klopries2024itf generates time series dataset by two interpretable methods. The first is the ITF-FM that operates the feature manipulation process based on the given dataset, The second is the ITF-GAN that is trained to generate indistinguishable data samples compared to the original data. ChatTS xie2024chatts comes up with the Attribute-Based Time Series Generator to get synthetic text-time series pairs for model training. LLMs are used to select proper features according to real-world settings for each attributes like trend, periodicity, local fluctuation and noise. The subsequent modules will generate the final time series based on the selection of the LLMs.
Graph Data Generation. To cope with the lack of large-scale graph structured datasets, a framework 10943225 to generate Large-Scale Graph Dataset is proposed. It generates structures and features separately by fitting the original graph, and aligns them to obtain the generated graph.  TKGGen presents the pipeline that generates temporal knowledge graph (TKG) datasets based on documents. It utilizes LLMs to generate timestamps for existing triplets, thereby obtaining the quadruples needed for constructing TKG.
3 LLM for Semi-Structured Data Analysis
Semi-structured data refers to data that are neither with strictly predefined schema like relational models nor raw data (e.g., plain text or images) abiteboul1997querying . Instead, they preserve partial organizational properties (e.g., tags, headers) and typically exhibit hierarchical or nested representation (e.g., County - Province - City in a nested JSON). This flexibility allows them to be represented in diverse formats, including web tables, spreadsheets, HTML, JSON, and XML. Broadly, such data can be categorized into markup languages and semi-structured tables, both of which pose challenges to downstream tasks due to their inherent structural complexity.
Takeaway In summary, semi-structured data remains difficult to handle due to irregular and nested structures.However, the strong semantic reasoning ability of LLMs shows great potential to overcome these challenges.Future research will focus on combining structure-aware methods with LLMs to achieve more robust and general solutions.
3.1 Markup Language
Markup languages (e.g., XML, JSON, HTML) are widely adopted for structuring and exchanging data. Nevertheless, research on integrating LLMs with markup languages remains relatively limited. Traditional methods for markup language understanding can be broadly classified into two categories. The first, Token Linearization, involves converting markup documents (e.g., HTML, JSON, Markdown) into plain text sequences and feeding them into PLMs herzig2020tapas . While straightforward, such methods are often too simplistic but hard to effectively capture the complexity of structured documents. To overcome this limitation, a second line of research, Tree/Graph-Based PLMs, seeks to explicitly encode structural information (e.g., DOM trees). However, owing to the limited generalization capacity and short context windows of PLMs, these models still struggle with issues such as the need for extensive pre-training, difficulties in handling large-scale data, and poor scalability.
Recently, LLM-based approaches have emerged as a promising alternative, offering stronger semantic reasoning and adaptability for handling the complexity of markup data. We conducted some research on this, and the work is shown in Figure 4.
Figure 4: LLM for Markup Language Overview
Markup Extraction. Markup Extraction refers to the task of extracting information from semi-structured data such as HTML and XML into the form of structured ones such as relational tables. Traditional approaches apply heuristic rules (e.g., find the content within the <h1><h1> or <title><title> tags wang2002detecting ) or data templates (e.g., extract the product name from <<h1 class=“product-name”>> crescenzi2001automatic ) to extract key information from markup information. However, these methods always face challenges such as template drift or high annotation costs. By leveraging the reasoning capabilities of LLMs, it becomes possible to directly extract and interpret hierarchical relationships, attributes, and nested structures from data, thus eliminating the reliance of the extraction task on index information such as tags or templates.
Evaporate arora2023language is a prototype system that leverages LLMs to derive structured views from semi-structured data. Given user-provided documents, it automatically generates structured representations of their contents. To support relational data extraction, Evaporate introduces the Evaporate-Code+ approach, which first employs LLMs to synthesize multiple candidate extraction functions. Specifically, the system generates functions through prompt engineering in two modes: one provides few-shot examples to the LLM and encourages the use of general-purpose Python functions, while the other specifies only the task requirements and directs the LLM to synthesize regular-expression–based solutions. Subsequently, a weak supervision–inspired aggregation mechanism is applied to assess and filter the candidate functions according to the quality of their outputs. Similarly, WebFormerwang2022webformer offers related functionality, but is tailored specifically to HTML by effectively incorporating its layout structure. Unlike Evaporate, WebFormer integrates website layout into transformer-based extraction by encoding DOM nodes as graph-attended HTML tokens and modeling text with efficient relative-position self-attention. Field tokens then guide heterogeneous cross-attention with HTML and text tokens, enabling accurate and scalable structured field extraction.
Markup Information Querying. Markup information querying aims to extract user-intended information from data encoded in markup languages. While conceptually analogous to SQL querying over relational databases, it differs in that markup files are typically irregular and loosely structured, thereby introducing distinct processing challenges. To address these challenges, existing research has primarily explored tree-based approaches and NoSQL-based methods.
$\bullet$ Tree-Based Approaches. XPath Agent li2024xpath adopts a two-stage pipeline for XPath query. In the Information Extraction (IE) stage, a lightweight LLM analyzes a simplified version of the webpage, where an HTML cleaning algorithm filters out irrelevant nodes and attributes, which finally generates a cleaned and compressed HTML. In the subsequent XPath Programming stage, a stronger LLM leverages the cleaned HTML and extracted information to construct XPath queries in a bottom-up manner. By progressively traversing from target nodes to parent elements and flexibly incorporating attributes such as class names and IDs.
$\bullet$ NoSQL-Based Approaches. SMART lu2025bridging is a Text-to-NoSQL agent that adopts a multi-step framework integrating small language models (SLMs), retrieval-augmented generation (RAG), and LLMs. Specifically, it first employs an SLM to generate predicted schemas and an initial NoSQL query solely based on the natural language query. Next, it performs RAG over the dataset by leveraging the generated schema, the preliminary NoSQL query, and the original query. The retrieved results are then incorporated as few-shot exemplars and combined with the intermediate outputs to guide the LLM in producing a refined NoSQL query. The resulting query can subsequently be applied to large-scale semi-structured data.
Markup Information Understanding. In contrast to querying markup languages, markup information understanding emphasizes leveraging LLMs to perform more complex operations on markup data (e.g., revision and error correction). To support such tasks, both end-to-end models and pipeline-based approaches have been developed for comprehensive understanding of markup files.
DOM-LM deng2022dom proposes a structure-aware Transformer for modeling HTML documents by jointly encoding textual content and the DOM tree structure. To handle large DOM trees, it partitions them into manageable subtrees and incorporates structure-aware positional embeddings (e.g., node depth, parent indices, and sibling indices) to capture hierarchical context. The model is further trained with a self-supervised masking strategy applied across both tokens and DOM nodes. In contrast, MarkupLM li2021markuplm extends token representations with four complementary perspectives. In addition to text embeddings for encoding document words, 1D positional embeddings for capturing horizontal order, and segment embeddings for distinguishing paragraph information, it introduces XPath embeddings. The method first decomposes an XPath into hierarchical layers consisting of tag names and subscripts. For each layer, the embeddings of the tag and its corresponding subscript are combined to form a unit-level embedding. These embeddings across all layers are then concatenated and passed through a feed-forward network (FFN) to generate the final XPath representation. For pre-training, MarkupLM employs three markup-specific objectives: (1) Masked Markup Language Modeling (MMLM) to jointly capture textual and structural information, (2) Node Relationship Prediction (NRP) to learn hierarchical relations among DOM nodes, and (3) Title–Page Matching (TPM) to assess global document coherence.
WebLM xu2024hierarchical introduces a hierarchical multimodal pre-training framework for webpage understanding. Specifically designed for HTML, it integrates structural information, textual content, and webpage screenshots within a unified Transformer encoder. The framework first aligns visual features with their corresponding HTML nodes and then employs two novel training objectives: (1) Tree Structure Prediction (TSP), which predicts parent–child and sibling relations among HTML nodes to strengthen structural understanding, and (2) Visual Misalignment Detection (VMD), which perturbs the size of node-associated visual regions to improve robustness against layout variations.
Insights of Markup Language Analysis. Currently, most studies on markup language understanding continue to rely on pre-trained language models (PLMs) for two main reasons. (1) Controllability: PLMs offer a fully controllable environment, where model size, training data, and task-specific settings can be explicitly defined. In contrast, LLM-based APIs are typically closed-source and involve high costs for fine-tuning and inference. (2) Structural Adaptation: PLMs can be tailored with specialized tokenization and positional encodings to capture the hierarchical structure of HTML/XML deng2022dom ; xu2024hierarchical . However, general-purpose LLMs often depend on prompt engineering, resulting in redundant tokenization and excessively long prompts, while also struggling to jointly model semantic and structural information (e.g., converting an HTML table into plain text requires an auxiliary schema to record the layout of each cell).
However, given the strong semantic capabilities of LLMs, leveraging them for markup language understanding remains promising, provided that effective methods are developed to compensate for their limitations while maximizing their semantic reasoning power. To this end, we identify two key research insights, which can be categorized into information compression and tool integration.
$\bullet$ Information Compression. Information compressiong aims to address the challenges of encoding markup languages by providing LLMs with more concise representations. Traditional approaches, such as structural prompting or schema-aware prompting, often result in redundant rules and excessively long prompts. In contrast, segment-wise prompting partitions long documents into manageable segments and selectively filters useful paragraph information. For renderable languages such as HTML or LaTeX, vision–language models may further assist in file partitioning and alignment with encoded content. For instance, using the VLM to help with visual anchor extraction, and generating a tree-structure to encode the HTML or LaTeX document. Then, we can flexibly select the anchor information according to the task and give it to LLM, which also reduces prompt length and retains the necessary structure information.
$\bullet$ Tool Integration. Tool Integration for LLM is an insight inspired by LLM Agent. This approach is to enable LLM with abilities to change the outer environment. There are already existing methods such as DSL-Based Soft Tools, Retrieval-based Code Generation, and API Calling frameworks. However, those tools are usually a little rigid, like API Calling frameworks is a fixed mapping between natural language and predefined functions.So we need to explore more dynamic and adaptive approaches, such as integrating planning modules, learning-based tool selection, and even self-reflective reasoning loops.
3.2 Semi-Structured Tables
In this section, we will focus on semi-structured table processing, which refers to techniques for understanding, representing, and extracting information from semi-structured tables.
Compared to structured relational data, semi-structured tables exhibit more complex structural organization, characterized by merged cells, hierarchical headers, and nested tables. In order to give a relatively clear definition of semi-structured tables, we summarize the common characteristics and representation formats of these tables as follows for your reference.
(1) Features of Semi-Structured Tables: We observe that many semi-structured tables share a set of recurring properties, which can be summarized into five core characteristics, illustrated in Figure 5. (1) Wrong Index: the indices of rows or columns may be missing, ambiguous, or inconsistent suhara2022annotating . (2) Hierarchical Content: not all rows or columns adhere to the same schema yang2022tableformer (e.g., a “summary” cell may appear under the “ID” column without any logical connection between the index and the content). (3) Merged Cell: a single cell may span multiple rows or columns guo2022trust . (4) Flexible Header Orientation: the relative position of keys and values is not fixed (e.g., keys may appear to the left of or above their associated values, as observed in the “summary” and “Name” fields). (5) Inconsistent Content: cell values are highly flexible and may include words, numbers, or even entire sentences.
Figure 5: Example Characters of Semi-Structured Tables.
(2) Table Representation Forms: The representation of semi-structured table is varied. We have classified and analyzed common formats and analyzed their characteristics as much as possible for your reference. The first thing we consider is whether the table format can support the representation of semi-structured tables. If so, can it fully represent the special structure of all tables? So we put forward “Semi-Structured Table Support” and “Completeness” to explain these two issues. Besides, we also investigated the structural complexity of different tables and proposed “Structural Complexity”, “Schema Complexity” and “Overall Relative Token Length Level”. “Structural Complexity” considers the complexity of the representation of the external structure of the cell, while “Schema Complexity” considers the complexity of the schema corresponding to each cell in the table structure. And the “Overall Relative Token Length Level” describes the length of the token formed by each form when prompting. This takes into account the complexity of the structure and schema, and is divided into levels 1 to 5 for your reference.
Through our research, we can easily find that tables with simple structure or annotations are often difficult to represent complex semi-structured table structures.(e.g.,CSV, Markdown).At the same time, although some forms (e.g.,JSON) have simple structures, they can be represented by complex schemas.
This inherent complexity poses substantial challenges in aligning queries with both the content and the structure of the table in query answering tasks. The absence of efficient tools (e.g., the openpyxl library for spreadsheet manipulation) and robust representation formats (e.g., Excel or HTML) further exacerbates the difficulty of processing such data.
However, with the aid of LLMs, it becomes possible to directly reason over semi-structured tables, interpret their intricate layouts, and bridge the gap between natural language queries and heterogeneous table structures, thereby enabling more accurate and context-aware query answering.
Traditional table processing approaches like TAPAS herzig2020tapas mainly rely on lightweight models (e.g., BERTdevlin2019bert , RoBERTaliu2019roberta ) to analyze the table. These methods rely on specialized embedding mechanisms to capture the two-dimensional structure of tables, including rows, columns, and numerical order. To support tasks such as cell selection or aggregation, they usually require large-scale weakly supervised pre-training yang2022tableformer . However, due to the small number of parameters and weak NLP capabilities of lightweight models, this method cannot handle some tasks based on semi-structured tables with confusing layouts.
Leveraging the strong semantic understanding capabilities of LLMs, LLM-based approaches generally preprocess the semi-structured table and then perform table understanding or reasoning. We conducted some research on this, and the work is shown in Figure 6.
Figure 6: LLM for Markup Language Overview
$\bullet$
Table Representation. By converting semi-structured tables into structured representations, LLMs can better capture the underlying logical relationships within tables, thereby improving their effectiveness in downstream tasks. However, the diverse and complex layouts of semi-structured tables (e.g., hierarchical headers, merged cells, and nested subtables) pose significant challenges to this process. To address these issues, tree-based modeling approaches and model-driven methods have been proposed.
(1) Tree-based Table Modeling: , TUTA wang2021tuta introduces a tree-based encoding framework for semi-structured tables. The root node branches into two child nodes, representing the row and column indices, which are further expanded into hierarchical trees. Each table entry is modeled as a node whose parents are the corresponding row and column nodes. For example, in a personnel information form, the column tree may contain nodes such as “ID” or “Phone Number”, while the row tree represents individual names. ST-Raptor tang2025st proposes a HO-Tree structure to comprehensively encode semi-structured tables. The HO-Tree consists of a Meta-Tree, which captures both the structural and semantic information of table headers, and a Body-Tree, which represents the structure and content of the table body. To construct the HO-Tree, the framework leverages a VLM to detect meta-information, applies heuristic rules to segment tables into subtables, and then builds the HO-Tree using a depth-first search strategy.
(2) Model-based Structuring: TabFormer 10.1145/3726302.3730071 introduces a framework for converting semi-structured tables into relational data. Its chain-of-thought (CoT) module first prompts LLMs to perform step-by-step transformations using “soft operators,” which are table manipulation tools provided through prompts (e.g., split operator to devide the table into subtables if multiple groups of headers are detected). After preprocessing, TabFormer leverages SQL for table reconstruction. Specifically, it first predicts the table indices, then generates the CREATE statements to define the table schema, and finally inserts the extracted contents from the pre-processed tables into the reconstructed relational tables. Auto-Tables li2023autotablessynthesizingmultisteptransformations integrates OCR for structure recognition and applies domain-specific language (DSL) operators. These operators abstract common spreadsheet operations (e.g., melt, transpose), thereby enhancing generalization and adaptability in table conversion tasks. TradeSweep lee2024can , while closely related to Auto-Tables, differs in its approach. Instead of predefined DSL operators, it accepts user-specified transformation requirements and directly generates table conversion code by leveraging a code repository, offering greater flexibility in addressing diverse user needs.
$\bullet$
Table Prompting. Table Prompting means converting table data into LLM’s text input format. The traditional way of table prompting is more like table lineriazation, which simply flattens the contents of all cells into a single string. This crude method will lead to a series of problems such as the prompt sequence being too long and the loss of structured information. Therefore, when using LLM to parse tables, we need to compress the structure and content information of the table. This can not only process large-scale text, but also effectively ensure that the loss of structural information is within a controllable range.
SHEETCOMPRESSOR dong2024spreadsheetllm introduces an efficient framework for serializing large-scale spreadsheets for LLM prompting. Specifically, Structural Anchors identify heterogeneous rows and columns that provide key layout cues, pruning redundant homogeneous ones to yield a skeletonized spreadsheet. Inverted-Index Translation replaces row/column serialization with a lossless JSON dictionary that indexes non-empty cell texts and merges identical values, thereby reducing storage overhead. Finally, Data Format Aggregation clusters adjacent numerical cells of the same type or format, capturing distribution patterns without redundancy. HySem pp2024hysem presents a lightweight LLM pipeline tailored for local deployment. Its Context Optimizer rewrites HTML table tokens into an extremely compressed form, ensuring each cell is represented by only one or two tokens (e.g., “Theme 3: Factors affecting stuff response” is compressed to “Theme 3”). This method establishes a direct mapping between markup and content by minimizing cell representations.
$\bullet$
Table Querying. LLMs typically require an effective and comprehensive method to extract table contents, a process often referred to as table querying. While this process is conceptually similar to SQL, in this section we focus only on approaches that extract information directly from semi-structured data, rather than traditional SQL-based querying.
CoSdong2024spreadsheetllm is a downstream table query program based on SHEETCOMPRESSORdong2024spreadsheetllm , which is mentioned above.First, given a compressed spreadsheet representation and a natural language query, the model performs Table Identification and Boundary Detection, locating the precise subtable relevant to the task. Second, the extracted subtable and query are passed back to the LLM for response generation, enabling focused and context-aware reasoning.
In ST-Raptortang2025st , the table query is based on its HO-Tree. A complex question is first decomposed into sub-questions, each mapped to a sequence of Basic Operations of Ho Tree (e.g.,Retrieval, Condition). To ensure these operations can be reliably executed, the system introduces an Alignment Operation, grounding abstract query terms to actual table headers or values. During execution, the model further leverages Top-down and Bottom-up retrieval strategies to flexibly navigate the hierarchical structure of HO-Tree. For large-scale tables, Column-Type Annotation (discrete/continuous/non-structural) is employed to constrain the search space. Finally, a dual validation mechanism is applied: Forward Constraints check the consistency of intermediate results, while Backward Reasoning reconstructs the question from the generated answer to reduce hallucinations.
3.3 Semi-Structured Data for LLM
To facilitate research on table understanding and reasoning under challenging conditions, a number of datasets have been developed. Below, we introduce several representative datasets that are widely used for semi-structured table understanding.
TEMPTABQA gupta2023temptabqa contains 11,454 question–answer pairs centered on temporal queries while SPREADSHEETBENCH ma2024spreadsheet provides a challenging benchmark for spreadsheet manipulation, consisting of 912 questions derived from real-world scenarios. MiMoTable li2024mimotable targets reasoning across multiple sheets and files, with 1,719 queries distributed over 428 spreadsheets. Evaluation results on these benchmarks reveal a significant performance gap, ranging from 20% to 50% between state-of-the-art models and human performance, underscoring the need for further advances in this area. INFOTABS gupta2020infotabs is designed for the natural language inference task on semi-structured tables. It contains 23,738 premise–hypothesis pairs, where premises are Wikipedia infoboxes and hypotheses are short sentences. WikiTableQuestions pasupat2015compositional focuses on question answering over semi-structured HTML tables, comprising 22,033 question–answer pairs together with their associated tables. FeTaQA nan2022fetaqa is a free-form table question answering dataset containing 10k Wikipedia-based pairs of tables, questions, answers, and supporting table cells. Unlike extractive QA datasets, FeTaQA provides human-written free-form answers including entities and their high-level relations. For comparison, Table II summarizes the characteristics of different datasets. The table highlights structural complexity including total cell counts and the number of merged cells, as well as question–answering types, including: (1) Content Match, where the task is to locate the correct cells, (2) Numeric Computation, which requires performing arithmetic or aggregation operations, and (3) Semantic Awareness, which involves semantically related or reasoning-intensive tasks. For instance, INFOTABS contains a large number of natural language inference questions, whereas SPREADSHEETBENCH features long, complex natural language queries (often exceeding 50 words) that describe multi-step tasks.
TABLE II: Statistics of Semi-Structured Table Datasets.
Table Features Q&A Features
#-Cell #-Merged-Cell
Content
Match
Numeric
Computation
Semantic
Aware
TEMPTABQA 11–100 1–10 $\times$ $\times$
SPREADSHEETBENCH 101–1K 11–100
MiMoTable 11–100 1–10 $\times$ $\times$
INFOTABS 11–100 1–10 $\times$ $\times$
WikiTableQuestions 101–1K 11–100
FeTaQA 101–1K 1–10 $\times$ $\times$
4 LLM for Unstructured Data Analysis
Unstructured data refers to data that lacks explicit structure, as it does not adhere to a predefined schema bigdataanalysis . Additionally, it exhibits high variability in format, length, and modality, which further complicates its processing and analysis baviskar2021efficient ; wang2025aop .
Figure 7: Overview of Chart Data Analysis.
4.1 Chart
Takeaway: Chart is mainly a graph-based data presentation method to organize structured information into a form that is easy for human to observe. The future for LLM chart data analysis lies in designing a mechanism that mimics human attention for chart observation, whether it is end-to-end or pipeline-based reasoning.
A chart is a graphical method for representing data. Since humans generally process visual information more effectively than raw tabular data, charts visualize data in symbolic forms such as bars, pies, and lines. However, as charts are inherently designed for human perception, developing rule-based methods for comprehensive understanding remains difficult. Early approaches introduced an additional layer with a fixed vocabulary to simulate rules. For example, DVQA kafle2018dvqaunderstandingdatavisualizations employed LSTM networks to encode questions and CNNs to encode chart images, but the reliance on a fixed vocabulary limited generalization. Subsequent works such as SANDY kafle2018dvqaunderstandingdatavisualizations alleviate this issue by leveraging OCR ground truth to link chart elements with dictionary entries, though the accuracy of OCR recognition remained a critical bottleneck.
From a technical perspective, chart understanding poses unique challenges as it requires integrating visual perception with structured data extraction, numerical reasoning, and logical inference. As shown in Figure 7, tasks on charts show diverse features, requiring a multi-perspective understanding on the structure and content of charts. As a result, multiple tools are useful for chart understanding, each focusing on a specific side of the task. Research in this domain has primarily followed two paradigms: (1) pipeline-based methods, which decompose the task into modality conversion and semantic reasoning, and (2) end-to-end models, which directly map chart images and instructions to responses within a unified framework. In the following, we provide an overview of chart-related tasks and examine these two LLM-based technical paradigms.
Chart Captioning. Chart Captioning (or Summarization, Chart-to-Text) refers to a task on generating a descriptive caption for a given chart. The caption should contain key information or serve as a summary of the graph. Early studies primarily target common graphical features such as axes and legends. These features enable different charts can be summarized by successively mentioning all of the features, giving rise to planning-based pipelines for caption generation. For instance, Mittal et al mittal-etal-1998-describing propose feature-specific strategies for different types of graphs, while Reiter reiter-2007-architecture outline a four-stage data-to-text framework comprising (1) Signal Analysis to look for possible patterns and trends, (2) Data Interpretation to identify more complex messages, (3) Document Planning to decide which of the messages should be mentioned in the generated text, and (4) Microplanning and Realisation to create an actual text. Such pipeline-based approaches are stable and extensible in fixed scenarios but tend to be heavy and lack general applicability. Building on this line of work, ChartThinker liu2024chartthinker employs a large-scale dataset and a chain-of-thought strategy to enhance the logical coherence and accuracy of chart summarization. Leveraging LLMs, it first identifies the chart type and then iteratively generates reasoning steps to capture a holistic understanding of the chart.
At the same time, end-to-end solutions have also emerged. obeid-hoque-2020-chart introduces a transformer-based model that generates summaries containing data by taking both the underlying table and chart metadata as input. FigCaps-HF singh2025figcapshffiguretocaptiongenerativeframework fine-tunes a figure-captioning model using RLHF. Other approaches extend beyond captioning to jointly address captioning and question answering by encoding chart elements with a grounded text decoder. For instance, UniChart masry2023unichart is a pretrained model that incorporates both low-level tasks (e.g., extracting visual elements) and high-level tasks (e.g., chart interpretation) into its training objectives, enabling more comprehensive chart understanding.
Chart Question Answering. Chart question answering (Chart QA) requires a model to answer questions related to the content of a chart. Compared to the task of captioning a chart, QA also has to grab the key information of a chart. However, the task requires a more flexible understanding of the charts, since a question may focus on subtle and imperceptible features of the charts. As mentioned in ChartQAmasry2022chartqabenchmarkquestionanswering , the questions might involve both visual and logical reasoning on charts.
Researches relevant to chart QA explores specific aspects of Chart QA task. wu2024chartinsights addresses the underexplored area of low-level ChartQA tasks (e.g., identifying correlations) by introducing the ChartInsights dataset and a tailored “Chain-of-Charts” prompt strategy. For the explanability of answering, Charts-of-Thought das2025chartsofthoughtenhancingllmvisualization designed prompts for altogether four steps in data processing: extraction, sorting, verification and question analysis. For multimodal alignment, ChartMoE xu2025chartmoemixturediverselyaligned used Mixture of Expert architecture to bridge the modality gap, as well as building ChartMoE-Align, a dataset with nearly 1 million chart-table-JSON-code quadruples to conduct alignment tasks.
Besides, some researches noticed the few-shot ability of LLM and designed relevant end-to-end solutions. ChartGemma masry2024chartgemmavisualinstructiontuningchart , for example, is trained on instruction-tuning data generated directly from chart images, thus exploiting both high-level and low-level information of charts. mPLUG-Owlye2024mplugowlmodularizationempowerslarge is a general-purposed model that also fits the problem. Researchers design a two-stage method to align visual information with text: In the first stage, visual knowledge and abstractor modules are trained; In the second stage, LLM and abstractor modules are fine-tuned using LoRA while freezing visual knowlegde model.
Chart-to-Code. Chart-to-Code refers to generating executable code (e.g., in matplotlib) that can be rendered into chart images, thereby transforming charts into editable forms. Current solutions are primarily end-to-end and rely on VLMs. ChartMimic yang2025chartmimicevaluatinglmmscrossmodal and ChartMoE xu2025chartmoemixturediverselyaligned provide large-scale benchmarks with extensive chart–code pairs to support this task.
Text2Chart31 zadeh2025text2chart31instructiontuningchart introduces a reinforcement learning–based instruction tuning approach for chart generation without requiring human feedback. The models are trained under two phases: supervised fine-tuning followed by reinforcement learning with tailored rewards to enhance performance. Chen et al. chen2025breakingsftplateaumultimodal employ multimodal structured reinforcement learning to surpass the performance plateau of supervised fine-tuning in chart-to-code tasks. The study leverages a multi-granularity structured reward system that uses multimodal textual and visual feedback to guide the RL training. The rewards are rule-based at textual level and model-based at the visual level.
LLM-Based Data Synthesis. A primary challenge in Chart QA is the scarcity of high-quality, multimodal instruction tuning datasets. To address this, several works have focused on data synthesis. Using GPT-4, ChartLlama han2023chartllama introduces a sophisticated pipeline with three phases successively: Chart data generation, chart figure generation, and instruction data(code, QA pairs, narrative paragraphs, etc.) generation. The pipeline enables the dataset to cover a wide array of chart types and tasks including QA, summarization and even chart editing. Similarly, ChartBench xu2024chartbenchbenchmarkcomplexvisual adopts both desensitized and GPT-generated data. For QA diversity, the work provides GPT with 200 different question templates, with human checks to ensure their correctness. huang2025evochart presents EvoChart. It is a self-training method with three phases: Compositional chart generation, chart evaluation and refinement, and QA pair generation and training. The three phases will run cyclically until the completion of data synthesis.
4.2 Video
Takeaway: Through reformulating videos as structured token sequences that contains temporal information as well as visual evidence, LLM can track and understand specific events in object level, and perform downstream tasks
Video is inherently a sequence of frames whose spatial content evolves along the temporal dimension. Effective analysis therefore requires models to jointly capture spatial semantics and temporal dynamics wang2024grounded . Traditional approaches typically rely on computer vision backbones to extract frame-level representations, followed by aggregation mechanisms such as pooling, token merging, or attention to model temporal dependencies. While these methods provide reliable performance, the architectures are often manually engineered while the process is annotation-intensive and computationally inefficient for long video sequences, frequently sacrificing fine-grained temporal cues during compression.
Recent advances employ LLMs to enhance video understanding within multimodal frameworks. By representing videos as structured token sequences rather than isolated frames, LLM-based paradigms enable more coherent alignment between visual evidence and textual instructions. This shift addresses three major limitations of previous work: (1) computational inefficiency mitigated by adaptive frame selection and token reduction strategies; (2) loss of visual fidelity alleviated through selective filtering of redundant or irrelevant frames; and (3) weak temporal grounding improved by incorporating modules that explicitly encode temporal order and event duration shen2024tempme ; wang2024grounded . Rather than discarding existing vision encoders, lightweight adapters or temporal reasoning components are integrated to extend their capabilities while preserving efficiency.
Figure 8 illustrates the general pipeline for LLM-aided video analysis, where existing work primarily focuses on optimizing video preprocessing (e.g., frame extraction, temporal sampling, and feature extraction) and semantic understanding (e.g., temporal modeling, multimodal fusion, and LLM integration).
Figure 8: Overview of Video Data Analysis.
Temporally-Anchored. Existing video-language models often struggle with precise temporal localization, as it requires detecting sparse and transient signals within long and redundant video sequences. This involves (1) distinguishing relevant information from irrelevant frames (e.g., the first appearance or disappearance of a target object, the exact frames of a key gesture, the onset of a sound event) shen2024tempme , (2) identifying event boundaries (e.g., action start/end points, scene cuts, speech or audio segment boundaries), and (3) preserving semantic coherence over time (e.g., preventing association of an object’s attribute at $t_{1}$ with an unrelated action at $t_{5}$ while minimizing information loss and computational overhead) chen2024timemarker . Meanwhile, existing models face difficulties in handling videos of varying lengths, where fixed input sizes waste computation on redundant segments in short videos while uniform sampling misses critical events in longer ones.
To address the above challenges, Temporally-Anchored methods explicitly align model predictions or reasoning with specific time points or intervals in a video. Concretely, these methods combine dynamic sampling (Adaptively select frames based on visual or semantic importance instead of uniformly sampling frames. For example, chen2024timemarker dynamically adjusts the sampling interval by monitoring motion intensity, allocating denser samples near action boundaries while sparsifying redundant segments.), temporal positional encodings (Models inject explicit time-aware representations into the token sequence. For instance, wang2024grounded augments frame embeddings with temporal position codes, enabling the model to reason about relative event order and duration.), and flexible tokenization (Including training-free similarity-based merging shen2024tempme , which groups redundant tokens based on feature similarity, and learnable saliency-driven merging lee2024video , which leverages gradient-based saliency to adaptively retain informative tokens.)
Specifically, TimeMarker chen2024timemarker , a versatile Video-LLM based on LLaMA3-8B, is designed for high-quality video-grounded dialogue with temporal localization. The model introduces separator tokens when the scene changes to represent physical time and temporal separator tokens (like second 5.) to enhance temporal awareness. In addition, it incorporates an AnyLength mechanism for dynamic frame sampling and adaptive token merging, which adjusts the frames-per-second (FPS) and merges visual tokens to match the LLM’s context window and GPU budget. Similarly, Grounded-VideoLLM wang2024grounded augments the architecture with a temporal stream and discrete time tokens to enable explicit timestamp reasoning, trained through progressive video–language localization tasks. Seq2Time deng2025seq2time further extends this direction by adopting sequential knowledge transfer which learns temporal grounding in a self-supervised manner by converting ordered image or clip sequences into time-annotated training signals. Specifically, researchers use a 10-second clip from Kinetics-700 to serve as the fundamental unit. Multiple such clips are then concatenated into a sequence to simulate the progression of events in a longer video. Each clip is associated with a caption and its positional index within the sequence. The model is trained to predict or align the sequential order of these clips, thereby internalizing temporal order even without explicit timestamp labels.
Instruction-Aware Relative Temporal Localization. In contrast to temporally-anchored approaches that output absolute timestamps or time-interval tokens (e.g., “at 00:02:15”), the instruction-aware relative temporal localization focuses on answering relative temporal queries (e.g., “what happened after the second bell?”).
Early efforts often integrate video foundation models (e.g., Conv3D/I3D-style models, two-stream models (models with architecture that process video stream through temporal stream and spatial stream separately) and transformer-based video encoders) with text-only LLMs, but such systems often required heavy pretraining and complex multi-module designs.
RED-VILLM huang2024image demonstrates that existing image-based LLMs can be augmented with lightweight temporal adapters that align instruction semantics with temporal video features, enabling efficient instruction-conditioned reasoning without retraining a full video model. IVA-LLM li2024llms integrates a spatial feature interactor and a temporal frame selector within LLM blocks. When a query is asked by the user, it first utilizes the temporal frame selector to select the most relevant frames from the video based the user’s query. Once the temporal frame selector has identified the key frames,the spatial feature interactor allows the model to focus on specific visual regions or objects within a frame. Such architectures enhance efficiency and visual fidelity under token aggregation while maintaining strong temporal grounding aligned with natural language instructions.
Video Emotional Analysis. Video emotional analysis seeks to infer affective states, social dynamics, and well-being indicators from multimodal cues. Understanding and promoting individual well-being is essential for staff health and long-term successmuller2024predicting . Current tools for assessing emotional changes rely on time-consuming surveys and questionnaires, which fails to find the subtle, context-dependent nature of emotion and the need for synchronized processing of visual, auditory, and textual signals.
In order to provide the real-time feedback needed to raise awareness and change individual behavior, PERMA muller2024predicting aims to predict individual well-being during teamwork interactions. Specifically, they first record panoramic video using cameras. For each individual in the video, they compute nonverbal behavior features such as facial expression, head pose, 3D gaze direction, and environmental parameters like scene brightness. These are organized into time series and summarized into 125 statistical features. Finally, these features are aligned with self-reported PERMA survey responses, and machine learning models are trained to predict each PERMA pillar—treated as a classification or regression tasks to distinguish discrete well-being states Another system jadhav2023ai evaluates candidates during interviews by fusing face detection with HaarCascade classifiers(a face detection algorithm that identifies faces within an input image and provide the coordinates of each detected face adepu2020interviewee ), extracting audio features via Mel-frequency cepstral coefficients jadhav2023ai , and modeling visual and acoustic sequences using Inception-V3 and LSTM encoders. Textual data are pre-processed through standard NLP pipelines (tokenization, stop-word removal, possibly sentiment features) before integration, allowing the model to generate joint predictions of personality traits and performance from frame-level and clip-level inputs.
Object Detection. Object detection localizes and classifies object instances in images by assigning bounding boxes and category labels. In video analysis, this task is further extended to ensure temporal consistency and to track specific objects across frames.
Recent methods introduce architectures capable of isolating region-level tokens and maintaining object-specific trajectories over time. To capture fine-grained spatial and temporal details that are often ignored by former models , VideoRefer yuan2025videorefer employs a spatial–temporal object encoder that couples a spatial token extractor with adaptive temporal token merging, producing precise region representations for single-frame and multi-frame reasoning. To enhance video analysis precision and efficiency de2024video , which employs YOLO-V8 to detect objects and segment event-relevant regions, while Gemini-2.5-flash serves as a multimodal language backbone to transform these visual detections into structured textual reports. The framework introduces an incident–context fusion pipeline that aggregates temporal object trajectories and scene cues before passing them into Gemini, enabling the model to generate summaries that retain both fine-grained entity information and broader situational context. By coupling high-precision, anchor-free detection with multimodal reasoning, the system reduces dependence on exhaustive manual review and ensures that critical events are captured without sacrificing efficiency. sengonul2025abnormal offers a pipeline for abnormal event detection in surveillance that spans three stages: (1) preprocessing through background subtraction and noise reduction, (2) feature learning via an LSTM autoencoder modeling normal behavior trajectories, and (3) anomaly detection using distribution-based regularity scoring enhanced by local minima analysis for confinement of anomalous instances while achieving high AUCs on CUHK Avenue and UMN datasets.
Gesture and Behavior Detection. Gesture and behavior detection aims to identify fine-grained body movements (e.g., hand gestures, nods) as well as higher-level behaviors (e.g., approach/avoidance, turn-taking), typically leveraging pose estimation, temporal classifiers, and interaction models.
Recently, MLLMs have been explored as tools for feature extraction in gesture and behavior detection. To explore how multimodal large language models (MLLMs) can be utilized as part of the multimodal learning analytics (MMLA) process,  whitehead2025utilizing employed MLLMs to automatically derive postural behavior indicators in a collaborative physics learning task, demonstrating that generative pipelines can reduce reliance on handcrafted annotations while preserving analytical depth. In this framework, MLLMs processed video data to extract features related to postural behaviors, leveraging their capacity to interpret visual and spatial information. As a method to generate detailed quantitative measurements of 3D human behavior previously unattainable through manual efforts or 2D methods , HARMONI weng2025artificial integrates three-dimensional mesh estimation with spatial interaction modeling to analyze caregiver–child interactions in naturalistic environments. It first applies video tracking algorithms to isolate individual body trajectories, then employs a deep neural network to estimate 3D human mesh models for each frame, which are subsequently refined through optimization and temporal filtering.
Video Data for LLM. High-quality video–language datasets are critical for large language models (LLMs) to acquire temporal reasoning, multimodal alignment, and grounding capabilities beyond static images or text. Video inherently encodes temporal dynamics, causal relationships, and fine-grained nonverbal signals—all of which are essential for tasks such as action understanding, spatio-temporal reasoning, and multimodal planning which can boost the capability of VLMs. To enable LLMs to capture these capabilities, researchers have developed several complementary approaches to collect or generate video–language data. These approaches can be categorized into four parallel paradigms, each offering unique advantages in terms of annotation fidelity, scalability, and diversity.
$\bullet$ Human Annotation. Human-annotated video datasets offer the highest label fidelity and domain specificity, making them indispensable for tasks where nuanced human judgment is required.
Gesture and interaction datasets, such as whitehead2025utilizing and weng2025artificial incorporate manual verificationto ensure the accuracy and relevance of the labeled data. and curated gestures and interaction detection datasets. For emotion and behavior analysis, muller2024predicting and jadhav2023ai combine human annotation with multimodal feature extraction—covering facial, audio, and textual cues—to produce rich, fine-grained annotations. Object detection–oriented datasets like guo2024xs utilize staged annotation processes with multiple rounds of verification to guarantee consistency .
$\bullet$ Human-in-the-Loop / Assisted Annotation. Human-in-the-loop pipelines strike a balance between annotation quality and scalability by combining expert oversight with lightweight automation.This approach leverages lightweight automated modules to support annotators. Examples include the building process in TimeMarker’s temporal-separator tokens or the construction of VidThinker’s clip captions and key-frame selection modules. Additionally, annotation workflows often employ lightweight modules to assist human experts. For instance, temporal frame selection can be guided by models like VideoITG’s VidThinkerwang2025videoitg , an automated pipeline for clip-level captioning, instruction-aware retrieval, and fine-grained frame localization, as demonstrated in wang2025videoitg with the VideoITG-40K dataset. Such tools ensure fine-grained alignment across video frames by combining automated support with expert oversight.
$\bullet$ Automated Curation & Filtering. Fully automated curation pipelines prioritize scalability and efficiency, generating large datasets while employing built-in quality control mechanisms. These are end-to-end automated generation pipelines with built-in quality control, often implemented as multi-agent engines with dedicated Reviewer agents, or via detection to LLM flows (such as YOLO-V8 to Gemini). For example, yuan2025videorefer employs a multi-agent engine to automatically curate fine-grained spatio-temporal object-level video instruction data for the VideoRefer-700K dataset, where a dedicated “Reviewer” agent filters erroneous or mismatched samples. As an example of detection to LLM flows, de2024video implements a pipeline that begins with automated object detection using YOLO-V8, followed by the Gemini model to generate incident- and context-aware video summaries, thereby reducing the need for exhaustive manual review.
$\bullet$ Synthetic Generation. Synthetic data generation provides virtually unlimited scalability by creating aligned video–text pairs using generative models.Fully synthetic data generation produces aligned video–text pairs via text-to-video and domain-specific generators, offering high scalability. Models such as NUWA wu2022nuwa employ a unified 3D transformer encoder–decoder architecture capable of handling text, images, and videos in a single framework for multimodal conditional generation. NUWA-Infinity extends this to arbitrarily long or high-resolution sequences using a hierarchical autoregressive “patch-over-patch” strategy with context caching and position control. Text2Video-Zero khachatryan2023text2video demonstrates zero-shot text-to-video generation using Stable Diffusion without additional training, enriching latent codes with motion dynamics and reprogramming frame-level self-attention via cross-frame attention anchored to the first frame for temporally consistent, high-quality videos. Align Your Latents blattmann2023align converts an image-only latent diffusion model into a video generator by adding a temporal dimension to the latent space and fine-tuning only temporal components, enabling high-resolution, temporally coherent generation. Domain-specific generators such as SadTalker zhang2023sadtalker and DreamTalk ma2023dreamtalk synthesize talking-head videos from audio, while choreography-guided systems wang2024disco produce human dance sequences from pose trajectories. High-fidelity text-to-video frameworks, including ho2022imagen and Make-A-Video singer2022make , accelerate large-scale creation of scene-specific clips by extending powerful text-to-image backbones with spatial–temporal modules, substantially reducing the need for manual annotation. Collectively, these synthetic methods provide a scalable and automated alternative to traditional pipelines, delivering richly diverse, automatically aligned training data that accelerates the development of multimodal LLMs.
4.3 Documents
Takeaway: Effective document understanding has shifted from unimodal text processing to sophisticated multimodal architectures that integrate textual, spatial, and visual information. Retrieval-Augmented Generation (RAG) and synthetic data are becoming crucial for enhancing factual accuracy and model robustness.
This section focuses on the challenges and techniques related to visually-rich document understanding, encompassing formats like PDFs, web pages, and scanned reports xu2020layoutlm ; appalaraju2021docformer . Unlike plain text, these documents present information with complex, non-sequential layouts, incorporating visual elements such as tables, lists, figures, and footnotes alongside standard text. The primary challenge lies in parsing this intricate combination of structural, spatial, and semantic information into a coherent, machine-readable representation (e.g., vectors) xu2021layoutlmv2 ; huang2022layoutlmv3 .
TABLE III: Challenges of Unstructured Document Analysis.
Document Type Key Challenges / Required Capabilities
Short-form Documents
(e.g., Invoices, web pages)
Factual QA, concise summarization, key-value extraction.
Long-form Documents
(e.g., Scientific papers, reports)
Integrative comprehension, cross-document synthesis, handling distractors.
Multilingual Corpora
(e.g., Parallel/comparable documents)
High-fidelity translation, cross-lingual QA and summarization.
Core Challenges in Document Understanding. Document understanding is complicated. As shown in Table III, for short-form documents (e.g., invoices, web pages), a key bottleneck is form understanding jaume2019funsd , which involves tasks like accurate factual question answering by locating specific information amidst layout noise, generating concise summaries, and performing key-value extraction from structured content. For long-form documents (e.g., scientific papers, technical reports), the challenges escalate to integrative comprehension and cross-document synthesis, along with a robust capability for handling distractors. Finally, handling multilingual corpora requires not only high-fidelity translation but also complex downstream tasks like cross-lingual QA or summarization.
To tackle these challenges, the research community has developed several distinct technical paradigms. We categorize them into three primary approaches: (1) architecture designs ranging from unimodal to multimodal designs, with particular emphasis on fusion strategies and cross-modal interactions, (2) retrieval-augmented methods for enhancing factual accuracy and context, and (3) synthetic data generation techniques to address data scarcity and improve model training.
TABLE IV: Taxonomy of Multimodal Document Understanding Architectures.
Fusion Stage Architecture Type Representative Models Technical Characteristics
Early Fusion Single-stream Transformer LayoutLM Concatenated input, unified processing
CNN + Transformer Traditional pipelines Visual CNN, text Transformer, early concat
Intermediate Fusion Dual-stream Transformer LayoutLMv2, DocFormer Parallel text/image streams, cross-attention
Cross-Attention Transformer VLCDoC Inter-modal attention mechanisms
Encoder-Decoder OmniParser, CREPE Shared encoder, task-specific decoders
Late Fusion Multi-head Networks DLAFormer Shared backbone, separate output heads
Pipeline Integration OCR+Layout systems Independent modules, result combination
Unified Processing Vision Transformer ViLT Patch-based, modality-agnostic processing
Architecture Design for Document Understanding. As shown in Table IV, the evolution of document understanding architectures reflects a fundamental shift from processing individual modalities in isolation to sophisticated multimodal integration. Early approaches treated text, layout, and visual elements as separate channels, which often led to suboptimal performance due to the loss of critical cross-modal relationships inherent in document structure. Unlike plain text processing, document understanding inherently requires the simultaneous consideration of textual content, spatial layout, and visual appearance. This realization has driven the field toward multimodal architectures that can capture these complex interdependencies.
$\bullet$ Transformer-Based Multimodal Architectures. The transformer architecture has become the dominant paradigm for multimodal document understanding. Foundational models like LayoutLM xu2020layoutlm pioneer the integration of text, layout, and visual information within a unified transformer framework. Building upon this, LayoutLMv2 xu2021layoutlmv2 introduces visual-textual attention mechanisms, while LayoutLMv3 huang2022layoutlmv3 advances unified text–image masking for more effective pre-training. DocFormer appalaraju2021docformer contributes multi-modal self-attention with separate attention heads for different modality pairs. Recent paradigm shifts emphasize efficiency and generalization. For example, DocLLM wang2024docllm represents layout-aware generative models that eliminate the need for raw visual encoders, significantly improving efficiency, whereas ViLT kim2021vilt takes this further by processing visual inputs in a convolution-free manner to concentrate computation on cross-modality interactions.
$\bullet$ Multimodal Fusion Strategies. The effectiveness of multimodal document understanding critically depends on how different modalities are integrated. Early fusion concatenates features at the input level; for example, LayoutLM xu2020layoutlm creates a single input representation for each token by combining textual embeddings with 1D sequential and 2D spatial position embeddings derived from bounding boxes. Intermediate fusion learns modality-specific representations separately before integrating them at deeper network layers, as exemplified by the dual-stream designs of LayoutLMv2 xu2021layoutlmv2 and DocFormer appalaraju2021docformer . By contrast, late fusion combines predictions from separately trained models; the shared-backbone, multi-head design of DLAFormer wang2024dlaformer can be viewed as a variant of this strategy.
$\bullet$ Cross-Modal Interaction Designs. Beyond fusion timing, cross-modal interaction mechanisms determine how effectively models capture fine-grained relationships. Attention-based designs such as VLCDoC bakkali2023vlcdoc adopt a two-step “align-then-integrate” mechanism: Inter-Modality Cross-Attention (InterMCA) first establishes semantic links between modalities (e.g., using text queries to attend to visual keys and values to obtain visually enhanced text representations) followed by Intra-Modality Self-Attention (IntraMSA) to fuse the enriched representations with each modality’s own context. Prompt-guided methods like VisFocus abramovich2024visfocus inject natural-language queries directly into the visual encoder, where cross-attention generates a weight map that highlights task-relevant regions. Coordinate-aware designs such as CREPE okamoto2024crepe further extend interaction by triggering a parallel coordinate-decoding head during autoregressive text generation, allowing the model to directly regress the bounding box of generated text when a special </ocr></ocr> token is emitted.
$\bullet$ Specialized Architectural Innovations. Recent work explores specialized designs to improve universality, efficiency, and reasoning capability. End-to-end unified frameworks like DLAFormer wang2024dlaformer jointly handle layout detection, role classification, and reading-order prediction as a single relation-prediction problem, taking a document image as input and outputting detected objects with class and relation information. To address the need for task-agnostic solutions, OmniParser wan2024omniparser proposes a universal framework that unifies outputs across tasks into three sequences (i.e., center points, polygonal shapes, and text contents) and decodes key–value pairs through a two-stage process that first predicts the center point of a value and then conditions on it to generate its boundary and text. Efficiency-oriented approaches such as Rationale Distillation hsieh2024rationale train a smaller model (e.g., Pix2Struct) to predict both final answers and intermediate reasoning steps of a larger model, optimizing a weighted combination of rationale and answer losses. Finally, multi-document reasoning systems like VisDoM lei2024visdommultimodal adopt a dual-pipeline RAG architecture, retrieving evidence separately from visual and textual sources and fusing them only when their consistency exceeds a predefined threshold.
$\bullet$ Robustness to Modality Variations. Real-world document understanding systems must remain effective when one or more modalities are missing or corrupted. To address this challenge, MissModal ma2023missmodal introduces a missing-modality adaptation strategy that aligns representations across different modality combinations. By minimizing the distributional distance between features derived from complete data and those obtained from incomplete inputs, the model enforces consistency within a shared semantic space. Complementarily, MMP huang2024mmp explores masked-modality learning by training a projection function to generate “pseudo” representations for absent modalities (e.g., vision) using information from the remaining ones (e.g., text and audio). This dynamic projection enables the model to reconstruct missing signals during inference, thereby preserving performance even under severe modality dropouts.
RAG-Based Document Understanding. To address challenges of factual accuracy and external knowledge, RAG has emerged as a cornerstone architecture (lewis2020retrieval, ). In document understanding, these systems can leverage different evidence modalities.
$\bullet$ Text-Based RAG. The most direct approach involves retrieving from the textual content of documents. More sophisticated methods like RAFT (zhang2024raft, ) focus on optimizing the retrieval and post-processing of text chunks.
$\bullet$ Image-Based RAG. Other systems leverage VLMs to perform retrieval directly on the visual content. M3DOCRAG (M3DOCRAG, ) uses vision-language models to generate embeddings for each document page and computes cross-modal similarity with the query directly on these visual embeddings, completely bypassing OCR. SV-RAG (SV-RAG, ) fine-tunes a VLM to first search for relevant page regions and then generate an answer from the retrieved visual evidence.
$\bullet$ Dual-Modality RAG. Hybrid approaches combine the strengths of both modalities. VisDoM (lei2024visdommultimodal, ), for example, implements parallel visual and textual retrieval branches and synthesizes evidence from both to formulate the final answer, improving reasoning by training models to focus on relevant retrieved context while ignoring distractors (zhang2024raft, ; yan2024corrective, ).
Document Synthetic Generation for LLM. Beyond processing existing documents, a burgeoning area of research focuses on using generative models to synthesize them. For instance, pisaneschi2023automatic adopts transformer-based models to generate plausible scientific paper layouts. A particularly clever approach, PosterLlama seol2024posterllama , reformats the layout generation task as HTML code generation. Instead of outputting coordinates directly, it instructs an LLM to generate HTML and CSS code describing the document layout, leveraging the model’s inherent knowledge of web layout languages. To overcome one-shot generation limitations, LayoutCoT (shi2025layoutcot, ) uses Chain-of-Thought (CoT) prompting for iterative refinement, while VASCAR (zhang2024vascar, ) introduces a visual self-correction loop. Meanwhile, diffusion models like LayoutDM (inoue2023layoutdm, ) and LDGM (hui2023ldgm, ) offer strong controllability. The frontier is moving towards aligning outputs with human aesthetic judgments, as exemplified by AesthetiQ (patnaik2025aesthetiq, ), supported by new datasets like SciPostLayout (tanaka2024scipostlayout, ) and metrics like LTSim (otani2024ltsim, ).
4.4 Program Language Analysis
Takeaway: The core of LLM-based program analysis is advanced data synthesis techniques. By creating high-quality “code-task pairs” through methods like autoformalization, iterative refinement, and RAG, models can tackle complex tasks from automated theorem proving to repository-level code completion.
In this section we introduce LLMs for program analysis, focusing on code-task pairs. Unlike raw source code, this data explicitly links program artifacts to their semantic properties or desired outputs, forming the basis for sophisticated analysis tasks. Examples include code paired with vulnerability labels, formal proofs, or natural language summaries. We first survey the downstream applications that this data enables and then describe the core methodologies for synthesizing this specialized data. Table V summarizes the core applications and the synthesis techniques that enable them.
TABLE V: Core Techniques in LLM-Based Program Analysis.
Application Synthesis Methods Key Techniques
ATP Task-Specific Generation
Iterative Refinement
Autoformalization
Expert Iteration
Vulnerability Detection Task-Specific Generation
Retrieval Prompting
Analysis Augmentation
In-Context Learning (ICL)
Code Summarization Few-Shot Learning
RAG
Synthesized Data Pairs
Example Retrieval
Repo-Level Completion RAG Adaptive/Iterative Retrieval
Context Fusion
Automated Theorem Proving (ATP). As a frontier application, ATP requires models to generate formally verifiable mathematical proofs. Recent data synthesis pipelines combine three key techniques to reach state-of-the-art performance on challenging benchmarks. Autoformalization translates natural-language mathematical problems into formal, machine-verifiable statements in languages such as Lean4. For example, DeepSeek-Prover xin2024deepseek-prover automatically generates large-scale Lean4 proof data from mathematical competition problems. Building on this foundation, expert iteration repeatedly tasks a model to solve a wide range of problems, collects all successful solutions as new training data, and then re-trains the model to become a stronger expert in the next round. Finally, reinforcement learning with verifier feedback allows models to explore proof strategies by generating candidate proofs and receiving binary correctness signals from formal verifiers (e.g., the Lean4 compiler), enabling learning through trial and error. The combination of autoformalization, iterative self-improvement, and verifier-guided reinforcement has empowered modern systems to generate entire, correct proofs in formal languages (xin2024deepseek-prover, ; lin2025goedel-prover, ; kimina2025prover, ).
Program Vulnerability Detection. LLMs enhance software vulnerability detection by learning from rich program data and leveraging contextual reasoning. One direction focuses on program-analysis–based training, where models such as PDBERT (liu2024pdbert, ) are fine-tuned on code augmented with program analysis information—such as control and data dependencies—to enable fine-grained vulnerability classification. Related methods also incorporate decomposed control flow graphs (CFGs) to capture intra- and inter-path features for more precise vulnerability prediction zhang2023bulnerability . Complementarily, case-driven prompt engineering exploits the in-context learning capability of LLMs by retrieving similar code snippets along with their known vulnerabilities and feeding this context, together with a model-generated interpretation of the target code, back into the LLM to boost detection accuracy. Approaches like VUL-GPT liu2023gptvulnerability and other prompt-based systems zhou2024llmvulnerability exemplify this strategy, demonstrating notable improvements in identifying subtle security flaws.
Code Summarization. By leveraging few-shot learning on synthesized code-summary pairs, LLMs significantly outperform traditional Seq2Seq models. Success relies on effective retrieval of similar code examples, using methods from token overlap to BM25, and enhancing prompts with semantic information like data flow to achieve state-of-the-art results (geng2023llmfewshot, ; ahmed2024code, ; mao2024scla, ).
Repository-Level Code Completion. LLMs are increasingly used to provide context-aware code completions by integrating information from an entire code repository. This task has progressed through a series of RAG techniques designed to resolve the so-called “Context–Latency Conundrum”, the trade-off between richer context and higher inference time. Early systems such as RepoCoder (zhang2023repocoder, ) adopt an iterative retrieval-and-generation pipeline, alternating between retrieving relevant code snippets and generating completions conditioned on the retrieved context. To reduce latency, REPOFUSE (liang2024repofuse, ) introduces efficient context fusion, combining an “analogy context” from similar code and a “rationale context” from semantic relationships, and employing Rank Truncated Generation to condense this information. More advanced frameworks such as Repoformer (wu2024repoformer, ) pursue adaptive and selective retrieval, allowing the model to self-assess when retrieval is actually beneficial and to skip it when unnecessary, thereby improving both robustness and inference speed. Pushing the conceptual boundary further, FT2Ra (guo2024ft2ra, ) proposes fine-tuning–inspired retrieval, treating each retrieval as a gradient-like update in which the retrieved document acts as a “gradient” and a tunable “learning rate” determines how strongly the retrieved information influences the evolving code representation. Together, these strategies illustrate a steady evolution toward repository-scale code completion that balances contextual richness with real-time responsiveness.
Data Synthesis for Code-Task Pairs. The ability to generate high-quality, large-scale datasets of code–task pairs is a primary driver of progress in this field. Two broad developmental paths characterize the evolution of code synthesis models. The path of intrinsic structure embeds formal grammars directly into model architectures to guarantee syntactic correctness. For example, early approaches such as SVAE (johnson2016composing, ) and SD-VAE (dai2018syntax, ) constrained the decoding process using explicit grammar rules, ensuring valid code outputs but limiting scalability due to structural rigidity. In contrast, the path of emergent capability leverages the powerful generalization ability of modern large language models through increasingly sophisticated data strategies. This trajectory advanced from scaling instruction diversity (Self-Instruct (wang2022selfinstruct, )), to systematically increasing instruction complexity (Evol-Instruct (luo2023wizardcoder, )), and finally to harvesting real-world code diversity from open-source repositories (OSS-Instruct (wei2023magicoder, )). The widespread adoption of functional-correctness metrics such as pass@@k further catalyzed the development of genuine problem-solving abilities in these models.
$\bullet$ Core Synthesis Techniques. At the heart of data synthesis lies task-specific data generation, where initial datasets are created for targeted analysis tasks. A representative example is autoformalization in automated theorem proving, which translates natural-language mathematical problems into formal, machine-verifiable statements (xin2024deepseek-prover, ; lin2025goedel-prover, ). Such task-driven pipelines provide the foundational training data needed to bootstrap powerful code models and to capture specialized reasoning patterns.
$\bullet$ Iterative Refinement and Self-Correction. Beyond one-shot generation, modern pipelines increasingly rely on iterative improvement strategies to enhance dataset quality. Expert iteration repeatedly tasks a strong model to solve large sets of problems, collects all successful solutions as new training data, and retrains the model to create progressively stronger experts (xin2024deepseek-prover, ; lin2025goedel-prover, ). Complementing this, reinforcement learning with verifier feedback allows models to explore solution spaces by generating code, receiving binary correctness signals from external tools such as compilers, and refining their outputs through trial and error. In self-debugging variants (chen2023teaching, ), models go further by using execution feedback to explain errors and autonomously correct their own outputs, dramatically improving sample efficiency and final code quality.
4.5 3D Models
Takeaway: Reformulating 3D models using the 3D encoder and projector aligns the 3D space with the semantic space, enabling llms to understand spatial features. Using domain-specialized interpreters or cross-modal feature extractor can improve the efficiency of 3D-llms.
3D models are digital representations of objects, scenes, or structures embedded in three-dimensional Euclidean space wikipedia_3dmodel . They are defined by explicit geometric data (e.g., point clouds, meshes, voxels, implicit surfaces) and may be further enriched with semantic or physical attributes (e.g., textures, materials, part-level annotations). Serving as a fundamental medium for spatial perception, reasoning, and interaction, 3D models provide the basis for computational systems to support tasks such as scene understanding, navigation, planning, and content creation.
Traditional 3D model analysis involves using 3D modeling software (e.g. Blender, 3ds Max, Maya) and need human input for manipulation, annotation, and interpretation. As LLMs evolve, their integration with 3D spatial data enables unprecedented capabilities in understanding and interacting with physical environments. This section provides a comprehensive overview of methodologies that allow LLMs to process, interpret, and generate 3D data from point clouds to Neural Radiance Fields (NeRFs), and examines their applications in tasks such as scene understanding, captioning, question-answering, and dialogue. We further discuss the role of LLM-powered agents in spatial reasoning, planning, and navigation, highlighting their potential to bridge natural language with the inherent complexities of 3D spaces. The fusion of LLMs with 3D data primarily manifests in three key areas: 3D Language Fusion (see Figure 9), which aligns 3D geometry with language descriptions; 3D-Derived Task Enhancement (see Figure 10), which leverages 3D information to improve performance on downstream tasks; and Cross-Modal Capability Refinement (see Figure 11), which uses existed experience on models (like 2-D image models) to refine 3D model’s understanding and generation abilities across different modalities.
TABLE VI: Overview of 3D-Language Fusion and Related Techniques.
Core Technical Objective Sub-Objective Details
3D-Language Fusion Encoder-Projection Architecture Dedicated 3D encoder + projection adapter; processes raw 3D data, and projects them into LLM-consumable feature space
End-to-End Tokenization Direct tokenization of 3D features; end-to-end fusion with text; no explicit projection layer
3D-Derived Task Enhancement Geometric Reasoning Textual abstraction of geometric problems; symbolic reasoning without computation; avoids direct spatial data exposure
Textual Description Mediation Converts 3D structures into textual descriptions; uses text-only model training/inference; avoids raw 3D input
Multi-Agent Orchestration LLM orchestrates specialized models; text-based API interaction; 3D data handled internally
Cross-modal Capability Refinement Domain Adaptation Cross-modal feature alignment; domain adversarial + contrastive learning; bridges 2D–3D gap
Feature Enhancement Modules Plug-in enhancement modules; addresses 3D perception weaknesses; preserves base architecture
Figure 9: Overview of 3D Model Analysis with 3D Language Fusion.
3D-Language Fusion. Fusing three-dimensional data representations with the semantic space of natural language aims to establish a unified modality that enables models to seamlessly interpret, reason about, and generate grounded descriptions of physical 3D structures. This process faces several challenges, including the inherent sparsity and irregular nature of 3D data (like point clouds), the significant computational cost of processing large 3D models, and the difficulty of establishing a precise semantic alignment between geometric features and linguistic concepts. As shown in Figure 9, the process includes using a 3D-Encoder to capture spatial features and relationships at first,and then uses a projector to align 3D features with text space so that the 3D model features can be injected into LLM well.
$\bullet$ Fusion with Point Clouds. 3D-LLM hong20233d pioneered this by converting point clouds or scenes into multi-view renderings, extracting per-view visual tokens with a frozen vision–language encoder, and aggregating them into 3D-aware tokens that are fed to the LLM, which attaches 3D positional embeddings and outputs location tokens aligned with language. Scaling to direct point-cloud input, 3UR-LLM xiong20253ur proposes an end-to-end 3D multimodal large language model that ingests raw point clouds, fuses them with text instructions, and projects them into a compact set of vision tokens via a 3D Compressor, which utilizes a multi-layer Transformer architecture with learnable queries and cross-attention mechanisms to interlink visual and textual features, and a 3D Query Fusion mechanism,which selects high-confidence queries from the perception component based on objectness probability and integrates them with the original queries from the compressor.
$\bullet$ Domain-Specialized Fusions. For scientific domains, 3D-MoLM li2024towards equips an LLM with a dedicated 3D molecular encoder (typically a graph-based neural network that encodes atomic coordinates and conformations) and a learnable molecule-text projector (usually implemented as a lightweight MLP or transformer adapter ) to encode atomic coordinates and conformations into language-context embeddings. The model is trained on a 3D molecule-text dataset to learn cross-modal retrieval, captioning, and open-ended QA.
ProteinChatguo2023proteinchat is designed to enable ChatGPT-like conversational functionality centered on protein 3D structures. Its architecture follows a four-part blueprint: (i) a protein encoder that derives sequence and 3D embeddings; (ii) a Protein–Language Pretraining Transformer (PLP-former), a Transformer-based module trained with contrastive and self-supervised objectives to align protein and language representations; (iii) a projection adapter, typically a lightweight neural network trained jointly with the PLP-former while keeping the LLM frozen, which converts protein embeddings into prompts suitable for language generation; and (iv) a foundation LLM (Vicuna-13B) that generates language-grounded responses.
Building on the framework of guo2023proteinchat , ProtChatGPT wang2024protchatgpt extends ProteinChat by introducing a three-stage pipeline:3D feature extraction, protein–language pretraining, and instruction tuning—to further enhance grounded question answering and explanatory capabilities for protein structures. This enhancement comes from its integration of both sequence- and structure-based protein encoders, contrastive pretraining that better aligns protein representations with textual descriptions, and instruction tuning that adapts the LLM for domain-specific reasoning, leading to more accurate and interpretable protein-grounded responses.
Figure 10: Overview of 3D Model Analysis with 3D-Derived Task Enhancement.
3D-Derived Task Enhancement Unlike 3D-Language Fusion which focuses on aligning spatial features with textual semantics, 3D-derived task enhancement emphasizes leveraging structural representations to improve downstream tasks that inherently require geometric reasoning (e.g., spatial question answering) as shown in Figure 10. The central challenge lies in (1) achieving genuine structural understanding rather than relying on superficial correlations, and (2) ensuring scalability, since incorporating high-dimensional 3D signals into large models can quickly lead to prohibitive memory and computation costs.
To evaluate LLMs’ understanding of geometric structures , GeomRel wang2025large identifies failure modes,including inaccurate identification of geometric relations, lack of structured reasoning in large language models (LLMs) by constructing a benchmark that isolates geometric relationship identification as a core subtask. To address this challenge, it encodes problems with structured relation pools (e.g., lines, angles, congruences) and labels target relations directly to facilitate accurate identification. For reasoning abilities, GeomRel introduces the Geometry Chain-of-Thought (GeoCoT), a two-stage prompting pipeline. The first stage forces enumeration and verification of candidate relations with reverse-checks for consistency, and the second stage performs rigorous reasoning over these relations. Ablations show that explicit supervision of relation identification helps LLMs learn spatial-structural semantics rather than producing numerically correct but conceptually hollow answers.
For molecular representation learning, GPT-MolBERTa wang20253dsmiles replaces compact SMILES weininger1988smiles strings, which linearize molecular graphs into text sequences, with rich textual molecule descriptions. The model is pretrained on 326k molecules with natural language captions to learn embeddings that capture geometric cues and the biological co-evolution to predict the property of target protein.
At system levels, ProtChat huang2024protchat integrates specialized Protein Large Language Models (PLLMs) into practical workflows with a multi-agent orchestration framework. A high-capacity planner LLM (GPT-4) decomposes user instructions, dispatches domain-specific PLLMs (e.g., MASSA hu2023multimodal ), and aggregates their outputs for tasks such as protein property prediction and protein-drug interaction profiling. The system standardizes prompt-tool contracts, result formatting, uncertainty reporting, and error handling, enabling non-expert users to leverage 3D biological models through a single natural-language interface without writing complex scripts.
Figure 11: Overview of 3D Model Analysis with Cross-Modal Capability Refinement.
Cross-Modal Capability Refinement. As illustrated in Figure 11, cross-modal alignment seeks to establish a shared representational space that enables seamless interaction among heterogeneous modalities, such as 2D images, 3D assets, and natural language, by leveraging knowledge transferred from other modalities. However, several challenges arise in this process, including the semantic gap between low-level geometric features and high-level linguistic concepts, as well as differences in data formats (e.g., continuous 3D point clouds versus discrete textual tokens).
To handle the limitation of domain adaptation technique in learning discriminative features for target samples due to the lack of label information in target domain, Song et al. song2023self improve image-mesh retrieval by combining adversarial domain alignment with self-supervised contrastive learning on unlabeled 3D data. The objective integrates (1) supervised discriminative learning on labeled 2D images, (2) contrastive losses on multi-view 3D renderings that pull views of the same object together and push different objects apart, and (3) adversarial terms to minimize domain gaps between 2D and 3D embeddings. An entropy-aware memory bank maintains representative negative samples and reduces false negatives, producing highly discriminative 3D embeddings and boosting retrieval accuracy.
To equip multimodal LLMs with 3D perception from single images, LLMI3D yang2024llmi3d augments a pretrained MLLM using parameter-efficient fine-tuning (PEFT) and three spatial reasoning modules: (1) Spatial-Enhanced Local Feature Mining sharpens object boundary and micro-geometry features via multi-scale attention with positional cues, (2) 3D Query Token-Derived Info Decoding introduces explicit geometric-query tokens with regression heads for precise numeric output, and (3) Geometry Projection-based Reasoning corrects for focal-length and perspective distortion by projecting features through differentiable camera models. The result shows that carefully designed spatial modules with PEFT can make general MLLMs competitive for 3D geometric reasoning without full-model retraining.
TABLE VII: Comparison of 3D Generation Methods across different paradigms.
Method Year Paradigm Uses 2D Prior
Multi-view
Consistency
Speed
Editable
Output
Key Techniques
Fantasia3D 2023 Optimization-based Medium Slow (minutes) SDS + DMTet mesh + BRDF
SweetDreamer 2023 Optimization-based ✓(tuned) High Slow (minutes) SDS + Aligned Geometric Priors
RichDreamer 2024 Optimization-based ✓(multi-prior) High Slow (minutes) SDS + Normal/Depth/Albedo Priors
Zero-1-to-3 2023 Feed-forward + Reconstruction Medium Fast (seconds) Limited View-conditioned diffusion + NeuS
Hunyuan3D-1.0 2024 Feed-forward + Reconstruction High Fast (seconds) Limited Multi-view diffusion + Sparse-view reconstructor
CraftsMan3D 2024 Native 3D Generation High Fast (seconds) 3D latent diffusion
Direct3D 2024 Native 3D Generation High Fast (seconds) D3D-VAE + D3D-DiT
MeshAnything 2024 Native 3D Generation High Fast (seconds) VQ-VAE mesh codebook + transformer
LLaMA-Mesh 2024 Native 3D Generation High Fast (seconds) LLM autoregressive mesh tokens
3D Data for LLM. Large language models have shown remarkable capabilities in reasoning over text and 2D imagery, but extending them to 3D domains introduces unique challenges. Unlike text or images, 3D data inherently encodes geometry, spatial relations, and multi-view consistency, which makes it both richer in information and more difficult to model. Recent research has therefore explored different paradigms for leveraging or directly generating 3D representations conditioned on language, aiming to unlock applications such as text-to-3D asset creation, embodied AI, and interactive design. Table VII Existing methods can be broadly grouped into three categories: optimization-based approaches that adapt powerful 2D diffusion priors for 3D generation, feed-forward pipelines that prioritize speed and determinism, and emerging native 3D generative models that attempt to learn 3D distributions directly.
$\bullet$ Optimization-Based. This paradigm leverages pre-trained 2D diffusion models as supervisory signals to directly optimize a 3D representation (e.g., NeRF or mesh). A key technique here is Score Distillation Sampling (SDS), which transfers the generative prior of 2D diffusion models into the 3D domain. In details , SDS adds noise to images rendered from the current 3D representation,and then uses the diffusion model to predict the denoising direction conditioned on the text prompt, then distills this signal into a gradient that updates the 3D parameters. Over many iterations, the 3D asset is shaped so that its 2D projections from diverse viewpoints are judged as likely given the prompt.
The primary advantage of this approach is its high quality ceiling:it can synthesize novel, intricate textures and surfaces by tapping into the rich visual priors of 2D diffusion models . However, it suffers from two major drawbacks: slow per-instance optimization (often taking tens of minutes) and multi-view inconsistency. The latter arises because the 2D prior lacks an inherent understanding of 3D consistency, often providing conflicting guidance from different views and leading to artifacts like the “Janus head” problem.
To address such challenges, Fantasia3D chen2023fantasia3d tackled the issue of blurred geometry in volumetric representations by explicitly disentangling geometry from appearance. It represents geometry with an explicit DMTet mesh and models appearance via a spatially-varying BRDF material model. During training, it separately optimizes geometry using normal map renderings and appearance using shaded renders via an SDS loss, resulting in high-fidelity, relightable assets. To combat the multi-view inconsistency problem, SweetDreamer li2023sweetdreamer proposed innovating on the prior itself rather than just the optimizer. It fine-tunes a 2D diffusion model on Objaverse renderings to predict view-conditioned canonical coordinate maps, termed Aligned Geometric Priors. This embeds multi-view consistency directly into the diffusion model, leading to more stable SDS gradients and reduced failures. Building on this, RichDreamer qiu2024richdreamer pushed prior specialization even further. Instead of relying on a single RGB diffusion model, it trains geometry-specific priors: one for normal maps, one for depth, and another for albedo (to factor out lighting). Applying SDS with these geometry-specific priors provides richer, higher-frequency supervisory signals, dramatically improving multi-view detail and surface fidelity compared to using a standard RGB-based SDS loss.
$\bullet$ Feed-Forward Generation with Reconstruction. This approach completely bypasses the slow, iterative SDS optimization process. Instead, it breaks the problem into two distinct, feed-forward stages: first, a multi-view synthesis model generates a set of coherent novel views from a single input image, and second, a fast 3D reconstruction network consumes these generated views to produce a 3D mesh in a single forward pass.
The key advantage is speed: generation time is reduced from minutes to seconds, offering deterministic behavior that is crucial for practical production pipelines. The trade-off is a loss of flexibility. The final output quality is bounded by the capabilities of the reconstruction network, and errors from the synthesis stage can propagate and accumulate in the reconstruction stage, potentially limiting the diversity and optimality of results.
This line of work focuses on improving the quality and coherence of the synthesized views to enable robust reconstruction. Zero-1-to-3 liu2023zero laid a foundational stone by training a viewpoint-conditioned 2D diffusion model to generate novel views from a single image. These synthesized views, while not perfectly consistent, are sufficiently coherent to be fed into an off-the-shelf reconstruction network like NeuSwang2021neus to produce a 3D mesh, all within seconds. Hunyuan3D-1.0 yang2024hunyuan3d follows this two-stage template but scales it into a robust, practical pipeline. It employs a powerful multi-view diffusion model that can generate a set of coherent RGB views from a single image in approximately four seconds. A key innovation is its learned sparse-view reconstructor, which is specifically trained to ingest these imperfect, AI-generated views (along with the original image) and predict high-quality geometry in a single, fast forward pass. This “synthesis-then-reconstruct” design prioritizes speed and determinism for rapid asset creation.
$\bullet$ Native 3D Generation. This emerging paradigm seeks to move beyond reliance on 2D diffusion priors altogether. The goal is to train generative models directly on large datasets of 3D assets, allowing them to learn the underlying distribution of 3D geometry and appearance natively. This requires designing efficient neural representations for 3D data, such as compact latent spaces or discrete tokens.
The primary advantage is inherent 3D consistency, eliminating multi-view issues from the root. It also promises fast generation and native support for 3D editing tasks. The main disadvantage is its heavy dependence on large-scale, high-quality 3D datasets for training, which are scarce compared to 2D image data. The technical complexity and computational cost of building these architectures are also currently very high.
Research is branching into different sub-architectures for native generation. One branch focuses on 3D diffusion. CraftsMan3D li2024craftsman3d performs diffusion directly in a latent space that represents both mesh and volumetric properties, enabling the generation of high-fidelity and editable meshes.To capture intricate geometry distributions,Direct3D wu2024direct3d introduces an architecture comprising a D3D-VAE that encodes a 3D asset into a compact latent representation, and a D3D-DiT (Diffusion Transformer) that learns to generate these latents, completely bypassing 2D priors and SDS. Another branch explores 3D tokenization, reframing 3D generation as a language modeling task.To optimize mesh extraction methods to no longer rely on dense faces and focus on geometric features , MeshAnything chen2024meshanything learns a discrete codebook of mesh components using a VQ-VAE, allowing a transformer to then autoregressively decode a sequence of tokens into a high-quality mesh. LLaMA-Mesh wang2024llama takes this concept to its logical extreme by fine-tuning a large language model (LLM) to directly output the vertices and faces of a mesh as a sequence of text tokens, effectively making the mesh a first-class citizen in language-conditioned generation and enabling fascinating applications like conversational mesh editing.
5 Heterogeneous Data Analysis
Takeaway: Heterogeneous data analysis requires capturing both structural and semantic information across diverse modalities, a task that traditional human-driven approaches handle inefficiently. LLM- and VLM-powered agents now enable cross-modal alignment, retrieval, and reasoning through natural language, significantly reducing manual effort and supporting more intelligent, context-aware analysis.
Heterogeneous data refers to the combination of diverse data types (e.g., relational data, semi-structured tables, document images) that exist in different formats. Effective analysis of such data requires the ability to capture both structural and semantic information, as well as the interrelationships among different modalities. Traditional approaches to heterogeneous data analysis focus on the construction of data lake systems for storage and retrieval. In these systems, analysts are responsible for writing retrieval queries and employing human-in-the-loop methods to conduct analyses, which are often labor-intensive and prone to error. Recent advances in LLMs and MLLMs offer new opportunities to analyze heterogeneous data through natural language interactions, thereby reducing human intervention. Current research primarily focuses on modality alignment across diverse data types and the development of heterogeneous data analysis agents.
LLM for Modality Alignment. Modality alignment refers to the process of mapping heterogeneous data types into a shared representation space that preserves both structural and semantic information. The main challenge lies in bridging the semantic gap among modalities with inherently different formats and levels of abstraction, while minimizing information loss.
Unicorn fan2024unicorn addresses the fundamental challenge of modality alignment, that determines whether two data objects convey the same semantic meaning. It employs DeBERTa he2020deberta to encode serialized pairs (e.g., text, table columns, tuples, or knowledge graph entities), incorporates a Mixture-of-Experts (MoE) architecture for feature alignment, and leverages an MLP-based binary classifier to assess whether the two objects match. Symphony chen2023symphony introduces a novel approach from a natural language perspective. The framework offers question–answering capability tailored for heterogeneous data lakes encompassing text, databases, tables, and knowledge graph entities. It transforms data from diverse modalities into natural language summaries, which are subsequently encoded into vector representations. Questions are encoded within the same vector space, and cosine similarity is employed to measure their alignment with the data representations, thereby enabling the retrieval of relevant information for answer generation.
LLM for Heterogeneous Data Retrieval. Data retrieval focuses on extracting task-relevant or related information from underlying data sources. The core challenge in heterogeneous data settings lies in the identifying and understanding of relationships both within and across different modalities.
Early efforts primarily targeted text-based modalities. For example, LOTUS patel2024lotus introduces a declarative programming interface that supports operations such as top-k search over data columns and the construction of similarity indexes for individual data items. Similarly, Wang et al. wang2025towards further abstract the retrieval process as a sequence of high-level APIs under a SQL-like paradigm, enabling operations such as selection, projection, join, and aggregation. While effective, these approaches remain confined to text data. Extending beyond this limitation, CAESURA urban2023caesura incorporates visual modalities by embedding visual question answering models into retrieval operations, thereby enabling multimodal queries such as counting the number of images relevant to a given query.
Heterogeneous Data Analysis Agents. A heterogeneous data analysis agent is an intelligent system that leverages LLMs or VLMs to automatically interpret, retrieve, and reason over heterogeneous data through natural language interaction. The key challenge is to ensure accurate, context-aware analysis while reducing reliance on human intervention, particularly in scenarios involving complex cross-modal reasoning.
XMODE nooralahzadeh2024explainable assumes that heterogeneous metadata is organized within a database and leverages LLMs to decompose natural language queries into subtasks such as text-to-SQL generation and image analysis. For instance, given the query “Plot the number of paintings that depict war in each century”, XMODE generates an operation sequence comprising: (1) retrieving painting metadata from the database via SQL, (2) applying VLMs to determine whether each image depicts war, and (3) producing visualization code. In practice, however, heterogeneous data is often disorganized. To address this challenge, Wang et al. wang2024interactive encode data objects using modality-specific models (e.g., BERT for text, CLIP for images) to construct an index. They further introduce a vector weight learning model that adaptively adjusts modality weights to better capture their relative importance in similarity measurement wang2024must . The user query is used to retrieve relevant data through semantic similarity for final answer generation.
6 Challenges and Future Directions
6.1 Structured Data
Complex Analysis. Structured data encapsulates both semantic information (e.g., cell contents in tables) and structural information (e.g., key–value relationships in tables or edge connections in graphs). Effective analysis requires integrating these two dimensions. For simple tasks, LLMs can often derive answers directly from explicit semantic cues or straightforward structural patterns. However, complex tasks demand deeper comprehension and multi-hop reasoning across semantics and structure, which remain challenging for current LLMs.
When LLMs alone fall short in handling such tasks end-to-end, multi-agent collaboration emerges as a promising direction. By decomposing a complex task into simpler subtasks, each agent can specialize in a narrower scope. Through prompt engineering or fine-tuning, LLMs can be tailored for specific subtasks, while intermediate outputs enhance both performance and interpretability.
Open World Task Adaptation. Structured data spans diverse domains with varying specifications and task requirements. LLMs trained for general-purpose analysis may fail to meet domain-specific needs. For instance, analyzing daily forms differs substantially from rigorous analyses in business or climate applications. Moreover, heterogeneity in system specifications, such as SQL dialects, poses additional adaptation challenges.
To address these issues, domain adaptation through task-specific training or fine-tuning is essential. As manual annotation is often insufficient to support large-scale training, efficient automatic data generation methods warrant further exploration. In addition, techniques such as automatic query conversion and intermediate representations are crucial for bridging modality gaps, particularly between natural language and diverse SQL dialects.
6.2 Semi-Structured Data
Diverse Markup Language Formats. Markup languages encompass a wide spectrum, including HTML, XML, JSON, Markdown, LaTeX, and domain-specific standards such as Journal Article Tag Suite (JATS). However, most existing research remains concentrated on a few widely used formats, primarily including HTML, XML, and JSON which often derived from web-scraped data. This leaves domain-specific or less-studied markup languages, as well as proprietary formats in industry, underexplored.
Addressing this gap requires the creation of specialized benchmarks for languages such as JATS, LaTeX, and Markdown to evaluate tasks like structure-aware question answering and format conversion. Moreover, model architectures must evolve towards structure-aware tokenization and training paradigms that explicitly leverage the hierarchical organization of markup, rather than treating markup as plain text sequences.
Understanding Complex Structured Table Semi-structured tables often exhibit irregular relationships and intricate nested structures, which remain difficult for current LLMs to fully capture. While recent approaches attempt to encode table structures and leverage the semantic reasoning capacity of LLMs, these methods are usually tailored to specific table types (e.g., financial tables). It is quite challenging to handle diverse semi-structured tables with complex structures.
To address this, researchers have explored several directions. One line of work focuses on Structure-Aware Encodings, where the table’s hierarchical or graph-like organization (e.g., DOM trees for HTML tables or JSON schemas) is explicitly modeled and injected into LLMs. Another approach emphasizes Intermediate Representations, such as transforming semi-structured tables into canonical forms or domain-specific languages (DSLs) that better align with LLMs’ sequential reasoning ability. In addition, Retrieval-Augmented Methods leverage external knowledge bases or schema repositories to supplement missing semantics, enabling LLMs to ground their reasoning in richer contexts.
Large-Scale Table Comprehension. For both PLM- and LLM-based methods, efficiency and accuracy are constrained by token limitations and context window sizes. When processing large-scale tables, effectively compressing both structure and content while minimizing information loss remains an urgent challenge. Advances in hierarchical encoding, table summarization, and scalable retrieval-augmented mechanisms may provide promising directions for addressing this limitation.
6.3 Unstructured Data
High-Level Chart Understanding. While significant progress has been made in basic chart recognition and data extraction, critical higher-order tasks such as chart-based fact checking, detection of visual or statistical misrepresentations (e.g., truncated axes, inconsistent scales, or fabricated trends), and cross-modal chart data verification remain largely unaddressed. These tasks demand not only accurate perception of visual elements but also reasoning about data semantics, statistical validity, and alignment with external knowledge or textual assertions. The absence of standardized benchmarks and evaluation protocols for such tasks further hinders systematic progress.
We advocate for the development of comprehensive frameworks that support critical chart literacy—enabling systems to not just “read” charts, but to audit them. This includes creating datasets with annotated errors, factual inconsistencies, and deceptive visual practices, as well as designing models capable of joint visual-semantic reasoning to validate the integrity and truthfulness of the graph.
Inadequate Multimodal Reasoning for Diverse Chart Types. Charts are inherently multimodal artifacts that fuse structured visual encodings (e.g., glyphs, spatial arrangements, color mappings) with natural language (titles, axis labels, captions). Current models often treat chart understanding as a vision-only or simplified vision-language problem, failing to capture the nuanced interplay between graphical conventions, domain-specific semantics, and contextual information. This limitation becomes especially pronounced when handling complex or non-standard chart types (e.g., small multiples, Sankey diagrams, or interactive dashboards).
A holistic, end-to-end chart understanding pipeline that unifies low-level perception (e.g., element detection), mid-level parsing (e.g., data-value mapping), and high-level inference (e.g., trend interpretation or anomaly detection) is needed. Such a pipeline should be modular, chart-type-agnostic, and grounded in both visualization theory and real-world usage patterns, enabling robust performance across diverse downstream applications.
Computational Efficiency of Video Analysis. Video is long-form sequential data; fixed sampling or full-frame processing consumes large compute and wastes capacity. Compressing frames/tokens often sacrifices fine-grained temporal cues, hurting localization and detail recognition.
To address this, we can develop representations that preserve fine-grained visual detail while offering controllable context length (e.g., temporal separator tokens, dynamic sampling, learnable token merging, temporal positional encodings). Integrate temporal modeling and token-aggregation objectives into training to achieve a better trade-off between efficiency and fidelity.
Video Temporal Localization with Semantic Coherence. Precise temporal localization requires detecting sparse, transient cues (e.g., action start/end, sudden events) and maintaining semantic consistency across long sequences. Current methods still struggle with event-boundary detection, timestamp reasoning, and preventing erroneous cross-time associations.
To address this,we can improve fine-grained event detection and semantic continuity by incorporating temporally aware tokens and supervision objectives directly into LLM-friendly representations.
Video Multimodal Fusion and Scalability. Effectively time-aligning and fusing visual, audio, and textual streams is both modeling and engineering challenging; high-quality, fine-grained human annotation is costly and hard to scale. Domain-specific tasks (emotion, gesture, anomaly detection, etc.) often rely on specialized labels and bespoke pipelines, which limits generalization.
To address this,we can combine modular architectures (lightweight temporal adapters, instruction-aware selectors), multi-agent data curation/reviewer pipelines, and large-scale synthetic video–text generation to improve annotation efficiency, domain adaptation, and instruction-conditioned reasoning. Encourage task-centered cooperative agents that decompose complex video-understanding problems into simpler subtasks.
Generalizing Multimodal Understanding for Documents. While models are moving beyond OCR, a core challenge is generalizing deep multimodal understanding—the holistic integration of text, layout, and visuals—across diverse domains and languages. General-purpose models often struggle with the unique structural grammars of specialized documents, such as the dense layouts of scientific papers or non-linear paths in historical manuscripts. Similarly, true multilingual generalization is hampered by language-specific conventions, like right-to-left scripts, revealing a gap between claimed capabilities and robust real-world applicability.
A dual approach is needed. First, developing efficient, document-native architectures (e.g., DocLLM) that prioritize layout information over expensive visual encoding. Second, enhancing model versatility through layout-aware instruction tuning that teaches reasoning about structure at multiple granularities. This must be supported by creating large-scale, high-quality datasets covering a broader range of languages and professional domains to enable true generalization.
Efficiency in Large-Scale Document Processing. Processing large volumes of long-form documents, such as legal archives, presents a trade-off between computational efficiency and analytical depth. The limited context windows of most LLMs force chunking strategies that can sever long-range dependencies, while preserving fine-grained spatial details across thousands of pages remains a major technical hurdle.
Future work should focus on adaptive context management and hierarchical document representations. Combining these with progressive reasoning strategies, where a model first builds a coarse understanding before focusing on details, can create a better balance between computational cost and high-fidelity analytical accuracy.
Accuracy and Robustness in Program Vulnerability Detection. While LLMs show promise in software vulnerability detection, the accuracy of state-of-the-art methods remains unsatisfactory, with reported success rates around 67.6% for detection and only 20% for automated repair. The number of vulnerabilities continues to grow, with over 25,000 new CVEs reported in 2024 alone. Moreover, the performance of deep learning-based methods has been shown to degrade significantly on rigorously validated datasets and in the face of simple code modifications.
The development of novel frameworks that integrate Code Property Graphs (CPGs) with LLMs presents a promising path. Techniques like CPG-guided code slicing can reduce code size while preserving vulnerability-relevant context, leading to significant improvements in F1 scores for inter-procedural vulnerability detection.
Program Analysis Complexity and Scope Limitations. The inherent complexity of vulnerabilities poses a significant challenge, with inter-procedural flaws being substantially harder to detect than intra-procedural ones. LLMs often struggle with uncommon CWE types, and their performance degrades when vulnerabilities span multiple code units. Most current LLM-based solutions are limited to the function level, a narrow scope that questions their effectiveness for class-level or repository-level analysis.
Future research should focus on developing LLM architectures that support multi-level code analysis, scaling from the function level to project and enterprise levels. Integrating Retrieval-Augmented Generation (RAG) techniques is crucial for effectively handling the vast contextual information present in large-scale codebases.
Reliability in Code Generation and Repair. Although LLMs offer advantages in generality and interpretability—providing intermediate reasoning for their findings—they still face significant challenges in terms of trust and accuracy, particularly for automated code repair.
Establishing mechanisms for trust and collaboration with developers is essential. LLMs should be designed to signal uncertainty rather than providing potentially inaccurate outputs. Furthermore, developing multi-agent collaborative frameworks, with specialized agents for code analysis, vulnerability detection, and repair, could lead to more robust and reliable systems.
High Quality 3D Modality Representation. 3D data come as point clouds, meshes, voxels, implicit fields, and NeRFs, each with different sampling, topology, and attributes. Converting these formats into compact LLM-ready inputs forces trade-offs: either aggressive simplification that loses geometric detail, or token explosion that breaks context budgets. These inconsistencies cause downstream reasoning to see uneven geometry and appearance cues, complicating generalization across tasks and formats.
To address this,we can develop compact, trainable encoders that convert raw 3D inputs into a small number of informative tokens suited for LLMs. Examples include vector-quantized mesh tokens, learned point-feature aggregators, and geometry-aware positional encodings. Jointly train these encoders with lightweight adapters using multiview-consistency, reprojection, and geometric-accuracy losses so tokenization preserves topology and metric cues while respecting LLM context budgets.
Accurate 3D Geometric Grounding. LLMs lack native metric reasoning and do not handle coordinate frames, units, or precise spatial numerics reliably. The scarcity of richly labeled 3D corpora with exact numeric annotations hinders supervised learning for distances, poses, and geometric constraints. As a result, systems may give plausible qualitative descriptions but fail on quantitative queries, undermining performance in navigation, manipulation, simulation, and safety-critical use cases.
To address this,we can use domain adapters that encode coordinate conventions, units, and regression heads for distances/poses, trained with metric supervision from simulation/renderers and differentiable-rendering or physical-constraint losses; evaluate on numerically grounded benchmarks to ensure quantitative fidelity.
6.4 Heterogeneous Data.
Pluggable Modality Extention Ability. A critical challenge for heterogeneous data agents is supporting the ability to incorporate new and diverse data types without redesigning or retraining the entire system. Traditional architectures are often rigid, requiring substantial modifications to handle emerging modalities such as audio, video, or sensor data. Without flexible integration mechanisms, agents risk becoming obsolete when encountering new data formats, limiting their long-term applicability and scalability.
Future work should focus on designing modular and extensible frameworks that allow seamless addition of new modalities through independent encoders or adapters. Research directions include developing standardized interfaces for modality integration, adaptive weighting strategies to balance multiple modalities, and self-supervised alignment methods to ensure that newly added modalities are compatible with existing representation spaces. Such approaches will enable heterogeneous data agents to remain scalable, maintainable, and robust in the face of evolving and increasingly complex datasets.
Heterogeneous High-Level Data Analysis. A key challenge in heterogeneous data agents is performing high-level analysis that goes beyond simple retrieval or extraction to include reasoning, summarization, and complex cross-modal inference. Unlike low-level tasks, high-level analysis requires understanding interrelationships among multiple modalities, capturing context, and generating actionable insights or visualizations. Existing systems, such as XMODE nooralahzadeh2024explainable and wang2024interactive , demonstrate partial solutions through step-wise query decomposition or adaptive modality weighting, yet they still struggle with disorganized or large-scale heterogeneous datasets, limiting their ability to provide fully automated, context-aware analyses.
Future research should focus on designing agents capable of robust, end-to-end high-level reasoning across heterogeneous modalities. Potential directions include leveraging multimodal embeddings for joint inference, and combining symbolic reasoning with neural models to ensure interpretability and correctness. Additionally, adaptive task planning mechanisms could enable agents to dynamically decompose complex queries into subtasks and synthesize results across modalities, thereby enhancing their capability for comprehensive, high-level data analysis in real-world scenarios.
7 Conclusion
In this paper, we summarize the recent techniques on LLM/Agent for data analysis, organized from the perspective of diverse data types, including structured data, semi-structured data, unstructured data, and heterogeneous data. Five key design goals for intelligent data analysis agents are distilled from the technical evolution. We also outline several remaining research challenges and propose some insights and practical directions for advancing LLM/Agents-powered data analysis.